### v3

# Improving the Euclidean Diffusion Generation of Manifold Data by Mitigating Score Function Singularity

Abstract
Euclidean diffusion models have achieved remarkable success in generative modeling across diverse domains, and they have been extended to manifold cases in recent advances. Instead of explicitly utilizing the structure of special manifolds as studied in previous works, in this paper we investigate direct sampling of the Euclidean diffusion models for general manifold-structured data. We reveal the multiscale singularity of the score function in the ambient space, which hinders the accuracy of diffusion-generated samples. We then present an elaborate theoretical analysis of the singularity structure of the score function by decomposing it along the tangential and normal directions of the manifold. To mitigate the singularity and improve the sampling accuracy, we propose two novel methods: (1) Niso-DM, which reduces the scale discrepancies in the score function by utilizing a non-isotropic noise, and (2) Tango-DM, which trains only the tangential component of the score function using a tangential-only loss function. Numerical experiments demonstrate that our methods achieve superior performance on distributions over various manifolds with complex geometries.
1 Introduction
11footnotetext: Corresponding author: tieli@pku.edu.cn (T. Li), wei.zhang@fu-berlin.de (W. Zhang)
Diffusion models [35, 11, 36, 38] have demonstrated remarkable success in generative modeling in various domains, including image generation [12], audio synthesis [5], molecule generation [13], and other applications [48, 22]. These models operate through a forward process, which gradually perturbs data into noise, and a reverse process, which reconstructs the data from noise.
Beyond Euclidean space, many scientific fields involve data distributions constrained on Riemannian manifolds. Examples include geographical sciences [28] using spheres, protein structures [45] and robotic movements [34] with $\mathrm{SE}(3)$ and $\mathrm{SO}(3)$, lattice quantum chromodynamics [26] with $\mathrm{SU}(3)$, 3D computer graphics [14] with triangular meshes, and cell development [21] with the Poincaré disk. To model such data, recent works [8, 26, 17, 49] have extended diffusion models to Riemannian manifolds, achieving notable progress by leveraging manifold geometry. Nonetheless, these methods face challenges due to the computational complexity of simulating diffusion and obtaining accurate geometric information like the heat kernel.
On the other hand, a natural idea is to directly apply diffusion models designed for Euclidean space to manifold-structured data. However, this approach encounters a fundamental challenge due to the singularity of the score function. Prior studies [29, 27, 3] have identified that under the manifold hypothesis, the norm of the score function explodes as the time in the reverse process approaches zero. Recent works [27, 29, 4] have theoretically analyzed this singularity under the Ornstein-Uhlenbeck process, providing asymptotic bounds on the score function and introducing structural assumptions on the data distribution. In this paper, we further enhance these insights and propose methods for mitigating the singularity of the score function under the manifold hypothesis.
Our work. We consider a probability distribution on a known $d$-dimensional submanifold $\mathcal{M}$ embedded in $\mathbb{R}^{n}$, where $d<n$. Given samples from this distribution, our objective is to generate new samples by directly applying Euclidean diffusion models. Within the framework of Variance Exploding Stochastic Differential Equations (VESDE) with such a manifold setting, we theoretically demonstrate that the perturbed score function exhibits distinct scales along the tangential and normal components. To address this issue, we propose the following two methods to improve diffusion models (DM):
  1. 1.
    Niso-DM: Perturb data with non-isotropic noise by introducing additional noise along the normal direction during the forward diffusion process. This reduces the scale discrepancy in the score function between the tangential and normal directions, making the relaxed score function easier to approximate.
  2. 2.
    Tango-DM: Train only the tangential component of the score function using a tangential-only loss function. By bypassing the training of the normal component, this method overcomes the learning difficulties caused by the multiscale issue.
Our main contributions are as follows:
  • We present an elaborate theoretical analysis of the multiscale singularity structure of the score function by separating it along the tangential and normal directions of the manifold in the embedded Euclidean space. Furthermore, we investigate the relationship between the full-space score function and the original Riemannian score function, which has not been considered before.
  • We propose two novel methods, Niso-DM and Tango-DM, making the relaxed score function easier to approximate. We also give a theoretical analysis of Niso-DM. Empirically, our methods demonstrate superior performance on distributions across various complex and non-trivial manifolds. This is in sharp contrast with previous methods that can only handle manifolds with special analytical structure.
2 Backgrounds and preliminaries
2.1 Diffusion models
Diffusion models can be formulated using stochastic differential equations (SDEs) [38]. Specifically, the data perturbation process is modeled by the following forward SDE:
$\mathrm{d}X_{t}=f(X_{t},t)\mathrm{d}t+g(t)\mathrm{d}W_{t},$ (1)
where $f$ and $g$ are fixed drift and diffusion coefficients, respectively. By carefully selecting $f$ and $g$, the distribution of $X_{T}$, with the probability density function (pdf) $p_{T}$, can be well approximated by a Gaussian distribution $\mathcal{N}(0,\sigma_{T}^{2}I)$, where $\sigma_{T}^{2}$ represents the variance of the Gaussian distribution. The corresponding reverse-time SDE, which transports $p_{T}$ back to the initial pdf $p_{0}$, is given by:
$\mathrm{d}X_{t}=\left(f(X_{t},t)-g(t)^{2}\nabla_{x}\log p_{t}(X_{t})\right)\mathrm{d}t+g(t)\mathrm{d}\overline{W}_{t},$ (2)
where $\overline{W}_{t}$ is a standard Brownian motion in reverse time and $p_{t}$ is the time-dependent distribution driven by the stochastic process $X_{t}$.
Inspired by this formulation, diffusion models employ a neural network $s_{\theta}(x,t)$ to approximate the score function $\nabla_{x}\log p_{t}(x)$. Since $p_{t}(x)$ is analytically unknown, direct optimization of the quadratic loss:
$\mathcal{L}_{\text{quad}}(\theta)=\mathbb{E}_{t\sim\mathcal{U}(0,1)}[\lambda_{t}\ell_{\text{quad}}(t,\theta)],$ (3)
$\ell_{\text{quad}}(t,\theta)=\mathbb{E}_{x_{t}}\|s_{\theta}(x_{t},t)-\nabla_{x_{t}}\log p_{t}(x_{t})\|^{2},$ (4)
is intractable. Instead, score matching techniques such as denoising score matching [43] are commonly used, which learn the score with the loss
$\ell_{\text{dsm}}(t,\theta)=\mathbb{E}_{x,x_{t}}\|s_{\theta}(x_{t},t)-\nabla_{x_{t}}\log p_{t}(x_{t}|x)\|^{2},$ (5)
where $p_{t}(x_{t}|x)$ denotes the probability density of $X_{t}$ in (1) conditioned on the initial state $X_{0}=x$. To utilize (5), the following two specific SDEs introduced in [38] are often considered, in which cases the density $p_{t}(x_{t}|x)$ has a closed-form.
  1. 1.
    Variance Preserving SDE (VPSDE):
    $\mathrm{d}X_{t}=-\frac{1}{2}\beta(t)X_{t}\mathrm{d}t+\sqrt{\beta(t)}\mathrm{d}W_{t},$ (6)
    where $\beta(t)>0$.
  2. 2.
    Variance Exploding SDE (VESDE):
    $\mathrm{d}X_{t}=\sqrt{\frac{\mathrm{d}\sigma_{t}^{2}}{\mathrm{d}t}}\mathrm{d}W_{t},$ (7)
    where $\sigma_{t}$ is a predefined noise scale, commonly chosen as $\sigma_{t}=\sigma_{\min}(\sigma_{\max}/\sigma_{\min})^{t/T}$ for some $\sigma_{\max}>\sigma_{\min}>0$. In this case, $p_{t}(x_{t}|x)$ follows a Gaussian distribution $\mathcal{N}(x_{t}|x,\sigma_{t}^{2}I)$.
In this study, we choose VESDE for its advantageous properties. Under VESDE, the mean of the perturbed distribution remains unchanged. At any time, the perturbed distribution $p_{t}(x)$ can be viewed as points on the original manifold $\mathcal{M}$ with added noise. This simplifies theoretical analysis and avoids the complications introduced by the evolving manifold in VPSDE. In fact, VPSDE involves an evolving manifold over time, defined as $\mathcal{M}_{t}=\{e^{-\frac{1}{2}\int_{0}^{t}\beta(s)\mathrm{d}s}x\mid x\in\mathcal{M}\}$ as the mean with noise perturbation.
2.2 Riemannian manifolds and notation
In this paper, the manifold $\mathcal{M}$ is viewed as $\mathcal{M}=\{x\in\mathbb{R}^{n}|\xi(x)=0\}$ embedded in $\mathbb{R}^{n}$, $\xi:\mathbb{R}^{n}\rightarrow\mathbb{R}^{n-d}$ is a known function, and we assume that $\nabla\xi(x)\in\mathbb{R}^{n\times(n-d)}$ has full column rank for all $x\in\mathcal{M}$. The target distribution is formulated as $p_{0}(x)\mathrm{d}\sigma_{\mathcal{M}}(x)$, where $p_{0}(x)$ is an unknown density and $\mathrm{d}\sigma_{\mathcal{M}}(x)$ represents the volume form [32] of $\mathcal{M}$. The unit normal vectors at a point $x\in\mathcal{M}$ are given by $N(x)=\nabla\xi(x)\left(\nabla\xi(x)^{T}\nabla\xi(x)\right)^{-\frac{1}{2}}$. For $x\in\mathcal{M}$, the operator $\nabla_{x}^{\mathcal{M}}$ is defined as $\nabla_{x}^{\mathcal{M}}h(x)=P(x)\nabla_{x}h(x)$ for any smooth function $h$, where $P(x)$ is the projection matrix [32] onto the tangent space of $\mathcal{M}$ and is given by
$P(x)=I-N(x)N(x)^{T}=I-\nabla\xi(x)(\nabla\xi(x)^{T}\nabla\xi(x))^{-1}\nabla\xi(x)^{T}.$ (8)
3 Scale discrepancy of score functions
It is commonly understood that singularities arise when Euclidean diffusion models are directly applied to data with manifold structures. In this section, we examine the singularity of diffusion models under VESDE, following the settings in Section 2.2. By introducing Gaussian noise across the entire space, the perturbed data become distributed throughout $\mathbb{R}^{n}$, deviating from their original confinement to the $d$-dimensional submanifold $\mathcal{M}$. The perturbed distribution is given by
$p_{\sigma}(\tilde{x})=\int_{\mathcal{M}}p_{0}(x)p_{\sigma}(\tilde{x}|x)\mathrm{d}\sigma_{\mathcal{M}}(x)=(2\pi\sigma^{2})^{-\frac{n}{2}}\int_{\mathcal{M}}p_{0}(x)e^{-\frac{|x-\tilde{x}|^{2}}{2\sigma^{2}}}\mathrm{d}\sigma_{\mathcal{M}}(x),$ (9)
where $\tilde{x}$ is a variable in the embedded space $\mathbb{R}^{n}$, $p_{\sigma}(\tilde{x}|x)$ denotes the pdf $\mathcal{N}(x,\sigma^{2}I)$ of the isotropic perturbtion, and $p_{0}(x)$ is only defined on $\mathcal{M}$. Hereafter, the subscript of $p$ denotes the noise scale $\sigma$ instead of the time $t$.
[…]
Theorem 3.1.
[…]
  1. 1.
    For $\tilde{x}\not\in\mathcal{M}$, assume that $x^{*}\in\mathcal{M}$ is the unique minimizer of $\min_{x\in\mathcal{M}}\|x-\tilde{x}\|$, and that $\|x^{*}-\tilde{x}\|_{\infty}<\frac{1}{n^{2}M_{1}}$. Under these conditions, as $\sigma\to 0$, we have
    $\nabla_{\tilde{x}}\log p_{\sigma}(\tilde{x})=\frac{x^{*}-\tilde{x}}{\sigma^{2}}+O(1)\,.$ (10)
  2. 2.
    For $\tilde{x}\in\mathcal{M}$, as $\sigma\to 0$, we have
[…]
    where $\frac{\partial P_{\cdot j}}{\partial x_{j^{\prime}}}$ denotes the vector whose $i$th component is $\frac{\partial P_{ij}}{\partial x_{j^{\prime}}}$.
[…]
$\ell_{\text{quad}}=$ $\mathbb{E}_{\tilde{x}}\|s_{\theta}(\tilde{x},t)-\nabla_{\tilde{x}}\log p_{\sigma}(\tilde{x})\|^{2}$ (13)
$=$ $\mathbb{E}_{\tilde{x}}\|P(\tilde{x})s_{\theta}(\tilde{x},t)-P(\tilde{x})\nabla_{\tilde{x}}\log p_{\sigma}(\tilde{x})\|^{2}+\mathbb{E}_{\tilde{x}}\|P^{\perp}(\tilde{x})s_{\theta}(\tilde{x},t)-P^{\perp}(\tilde{x})\nabla_{\tilde{x}}\log p_{\sigma}(\tilde{x})\|^{2}$
$=$ $:\ell_{\text{quad}}^{\parallel}+\ell_{\text{quad}}^{\perp}.$
Using Eq. (10) and the facts that $P(x^{*})(x^{*}-\tilde{x})=0$, $P(\tilde{x})=P(x^{*})+O(\tilde{x}-x^{*})$ and $\mathbb{E}_{\tilde{x}|x}(x^{*}-\tilde{x})=O(\sigma)$, the two terms being approximated in (13) have scales of $O(1)$ and $O(1/\sigma)$, respectively, by the following two estimates:
$\mathbb{E}_{\tilde{x}|x}[P(\tilde{x})\nabla_{\tilde{x}}\log p_{\sigma}(\tilde{x})]=\mathbb{E}_{\tilde{x}|x}\left[\left(P(x^{*})+O(\tilde{x}-x^{*})\right)\Big(\frac{x^{*}-\tilde{x}}{\sigma^{2}}+O(1)\Big)\right]=O(1),$ (14)
$\mathbb{E}_{\tilde{x}|x}[P^{\perp}(\tilde{x})\nabla_{\tilde{x}}\log p_{\sigma}(\tilde{x})]=\mathbb{E}_{\tilde{x}|x}\left[\left(P^{\perp}(x^{*})+O(\tilde{x}-x^{*})\right)\Big(\frac{x^{*}-\tilde{x}}{\sigma^{2}}+O(1)\Big)\right]=O\Big(\frac{1}{\sigma}\Big),$ (15)
where $\mathbb{E}_{\tilde{x}|x}$ denotes the expectation with respect to $p_{\sigma}(\tilde{x}|x)$.
This multiscale singularity of the loss formulation poses significant challenges during training. In fact, the above analysis also applies to the loss (5), since it differs from (4) only by a constant that is independent of $\theta$. As a result, training with (5) inherently prioritizes fitting larger-scale features of the score, which are aligned with the normal component in our settings. This is because quadratic loss minimizes the average of squared errors, making it disproportionately sensitive to errors that occur in larger-scale directions. In the context of this work, the normal component pulls samples back onto the manifold, whereas the tangential component, which is more critical, captures the data distribution on the manifold. As a result, the model fails to adequately capture finer details of the data distribution, ultimately reducing the accuracy of the generated distribution. In other words, the trained model primarily captures the manifold itself rather than the distribution on it, leading to a phenomenon known as manifold overfitting [24].
To address these limitations, specific methods are required during the model training phase to ensure training stability and accurately capture the intrinsic data distribution on the manifold. Based on this perspective, we propose the following two methods. The first method reduces the scale discrepancy between the tangential and normal components by modifying the structure of the noise. The second bypasses the learning of the normal component of the score function and employs a dedicated projection operator to project samples back onto the manifold.
4 Methods
In this section, we propose two methods to address the singularity of the score function. The first method, referred to as Niso-DM, introduces additional noise along the normal direction to mitigate its dominance. The second method, called Tango-DM, focuses only on the training of the tangential component of the score function when the noise scale $\sigma_{t}$ is small.
4.1 Niso-DM: Perturb data with non-isotropic noise
[…]
Theorem 4.1.
[…]
$p_{\sigma}(\tilde{x})=(2\pi)^{-\frac{n}{2}}\int_{\mathcal{M}}p_{0}(x)(\det\Sigma_{\sigma}(x))^{-\frac{1}{2}}e^{-\frac{1}{2}(\tilde{x}-x)^{T}\Sigma_{\sigma}(x)^{-1}(\tilde{x}-x)}\mathrm{d}\sigma_{\mathcal{M}}(x),$ (16)
[…]
  1. 1.
    For $\tilde{x}\not\in\mathcal{M}$, assume that $x^{*}\in\mathcal{M}$ is the unique minimizer of $\min_{x\in\mathcal{M}}\|x-\tilde{x}\|$, and that $\|x^{*}-\tilde{x}\|_{\infty}<\min\{1,\frac{2}{n^{2}(4M_{1}+M_{2})}\}$. Under these conditions, as $\sigma\to 0$, we have
    $\nabla_{\tilde{x}}\log p_{\sigma}(\tilde{x})=\frac{x^{*}-\tilde{x}}{\sigma^{2\alpha}}\cdot\frac{1}{1+\sigma^{2-2\alpha}}+O(\sigma^{(1-2\alpha)\wedge 0}),$ (17)
    where the symbol $\wedge$ is defined as $a\wedge b=\min\{a,b\}$.
  2. 2.
    For $\tilde{x}\in\mathcal{M}$, as $\sigma\to 0$, we have
    $\nabla_{\tilde{x}}\log p_{\sigma}(\tilde{x})=\nabla^{\mathcal{M}}_{\tilde{x}}\log p_{0}(\tilde{x})+O(\sigma^{(2-2\alpha)\wedge 1}).$ (18)
Eq. (17) shows that, by introducing additional noise along the normal direction, the scale of the normal component is reduced from $O(1/\sigma^{2})$ to $O(1/\sigma^{2\alpha})$. Eq. (18) states that this approach does not affect the characterization of the distribution on the manifold (i.e., the tangential component). As the noise level parameter $\sigma_{t}$ approaches zero, the score function of the perturbed density $p_{\sigma_{t}}$ on the manifold converges to the Riemannian score function of $p_{0}$. Therefore, by employing non-isotropic noise, we can still generate samples that adhere to the original distribution of the dataset.
Noting that $\nabla_{x_{t}}\log p_{\sigma_{t}}(x_{t}|x)=-\Sigma_{\sigma_{t}}(x)^{-1}(x_{t}-x)$, the denoising score matching loss (5) can be written into the following form:
$\ell_{\text{Niso}}(t,\theta)=\mathbb{E}_{x,x_{t}}\|s_{\theta}(x_{t},t)+\Sigma_{\sigma_{t}}(x)^{-1}(x_{t}-x)\|^{2}.$ (19)
Under the settings in Section 2.2, $\Sigma_{\sigma_{t}}(x)^{-1}$ has a closed-form expression (See (49)). Besides, we set $\alpha_{t}=\log c_{\mathrm{niso}}/\log\sigma_{t}$ (i.e., $c_{\mathrm{niso}}=\sigma_{t}^{\alpha_{t}}$) in practice, where $c_{\mathrm{niso}}$ is a fixed constant.
4.2 Tango-DM: Learn only the tangential component of the score function
Recall that the projection matrix $P(x)$ in (8) is well-defined in the ambient space $\mathbb{R}^{n}$. Accordingly, we define $s^{\parallel}_{\theta}(x,t):=P(x)s_{\theta}(x,t)$ for $x\in\mathbb{R}^{n}$, which represents the tangential component of the parametrized vector field.
As discussed in Section 3 (see (13)), the loss function (4) and (5) can be decomposed into two parts, $\ell_{\text{quad}}^{\parallel}$ and $\ell_{\text{quad}}^{\perp}$, and the singularity issue is associated with the part $\ell_{\text{quad}}^{\perp}$. To address this, we propose training only the tangential component of the score function using the loss $\ell_{\text{quad}}^{\parallel}$ when the noise scale $\sigma_{t}$ is sufficiently small, thereby avoiding the singularity associated with $\ell_{\text{quad}}^{\perp}$. To compute $\ell_{\text{quad}}^{\parallel}$, we introduce the following Tango loss
$\ell_{\text{tango}}(t,\theta):=\mathbb{E}_{x,x_{t}}\|s^{\parallel}_{\theta}(x_{t},t)-P(x_{t})\nabla_{x_{t}}\log p_{\sigma_{t}}(x_{t}|x)\|^{2}.$ (20)
The optimal score network $s^{\parallel}_{\theta^{*}}(x,t)$ that minimizes (20) satisfies $s^{\parallel}_{\theta^{*}}(x,t)=P(x)\nabla_{x}\log p_{\sigma_{t}}(x)$ (See Appendix C.1 for the proof). This result ensures the validity of the loss function in (20).
When the noise scale is smaller than a pre-selected $c_{\mathrm{tango}}$ (i.e. $\sigma_{t}<c_{\mathrm{tango}}$), the tangential component of the score function $s^{\parallel}_{\theta}(x,t)$ is trained via the Tango loss $\ell_{\text{tango}}$. On the other hand, when $\sigma_{t}\geq c_{\mathrm{tango}}$, the singularity issue is less severe, allowing us to use the original denoising score matching loss (5) to train the entire score function $s_{\theta}(x,t)$, which includes information along the normal direction. This enables the score function to push points toward the vicinity of the manifold. The overall loss calculation is summarized in Algorithm 1 in the Appendix.
5 Generation
Figure 1: The average error of the tangential component of the learned score function in the xyx-y plane, along the $z$-axis and $t$-axis, with $\sigma_{\min}=0.01$. From top to bottom, the plots correspond to the vanilla algorithm (Iso-DM), our proposed Niso-DM and Tango-DM, all using the rescaling technique.
We propose two sampling methods to generate samples using the learned score function $s_{\theta}$. The first method, named Reverse SDE, is based on the standard generation process in $\mathbb{R}^{n}$, with an additional projection step to ensure that the samples lie on manifolds. The second method, named Annealing SDE, adopts a two-stage approach, where the first stage simulates the standard reverse SDE, and the second stage performs annealed Langevin dynamics constrained to manifolds. The details of these methods are provided in the sections below and are summarized in Algorithms 2 and 3 in the Appendix.
5.1 Reverse SDE
The first method follows the standard generation process, where samples are generated by simulating the reverse SDE (21) with $t$ decreasing from $T$ to 0, which is derived from (2) by setting the drift term $f$ to 0.
$\mathrm{d}X_{t}=-g(t)^{2}s_{\theta}(X_{t},t)\mathrm{d}t+g(t)\mathrm{d}\overline{W}_{t}$ (21)
However, the ending samples lie in the vicinity of the manifold rather than exactly on it. To address this, we add a final projection step to ensure that the generated points lie on the manifold.
In general, this method is not applicable to Tango-DM, which utilizes the Tango loss in (20). Since the normal component of the score function is not learned when $t$ is small, the score function $s_{\theta}(x,t)$ does not capture information in the normal direction. For Niso-DM, the final distribution of the reverse SDE is the same as the distribution of $x+c_{\mathrm{niso}}N(x)\epsilon_{2}$, where $x\sim p_{0}$ and $c_{\mathrm{niso}}N(x)\epsilon_{2}$ represents noise along the normal direction. By employing a projection onto the manifold in the subsequent step, the obtained samples can accurately approximate the target distribution.
5.2 Annealing SDE on manifolds
The second method adopts a two-stage sampling procedure with a predefined threshold $\tilde{\sigma}$. In the first stage ($\sigma_{t}\geq\tilde{\sigma}$), we simulate the reverse SDE (21), which provides a coarse approximation of the target distribution. In the second stage ($\sigma_{t}<\tilde{\sigma}$), we refine the samples by simulating annealed Langevin dynamics constrained to the manifold. Specifically, for a fixed $t$, we perform $n_{0}$ steps of the following Langevin dynamics:
$\mathrm{d}X_{s}=P(X_{s})s_{\theta}(X_{s},t)\mathrm{d}s+\sqrt{2}\mathrm{d}W_{s}^{\mathcal{M}}.$ (22)
After that, we decrease $t$ according to $t\leftarrow t-\Delta t$ and restart the simulation of SDE (22) using the final samples from the previous iteration as the initial states. This process is repeated until the desired level of refinement is achieved. i.e., $t=0$. This stage can be viewed as the manifold version of the Corrector algorithms in [38].
To analysis the proposed sampling method, we assume that the neural network $s_{\theta}$ has learned the optimal solution, i.e. $P(x)s_{\theta}(x,t)=\nabla_{x}^{\mathcal{M}}\log p_{\sigma_{t}}(x)$ for $x\in\mathcal{M}$. Under this assumption, the invariant distribution of the SDE (22) is $p_{\sigma_{t}}(x)\mathrm{d}\sigma_{\mathcal{M}}(x)$. According to (12) in Theorem 3.1 and (18) in Theorem 4.1, regardless of whether non-isotropic noise is used, $\nabla_{x}^{\mathcal{M}}\log p_{\sigma_{t}}(x)$ converges to $\nabla_{x}^{\mathcal{M}}\log p_{0}(x)$ as $t\to 0$. This demonstrates that, by employing annealed Langevin dynamics constrained to the manifold, we can effectively sample from the target distribution $p_{0}(x)\mathrm{d}\sigma_{\mathcal{M}}(x)$ on the manifold.
6 Experiments
In this section, we evaluate our methods by studying several representative problems on typical manifolds. Specifically, we consider four datasets on four different types of submanifolds: (1) a hyperplane in 3D space, (2) mesh manifolds in 3D space, (3) the special orthogonal group of order $10$, and (4) the level set of dihedral angles in the molecular system alanine dipeptide. Our code is available at: https://github.com/ZichenLiu1999/NisoTangoDM.
We perform experiments with the vanilla diffusion models, as well as our proposed Niso-DM and Tango-DM, denoted as Iso, Niso, and Tango, respectively. Moreover, we consider using neural networks to approximate normalized score functions, defined as $s_{\theta}(x,t)=\hat{s}_{\theta}(x,t)/w_{t}$, where $\hat{s}_{\theta}(x,t)$ is a neural network and $w_{t}$ is the scaling factor of the optimal score function. This rescaling technique improves numerical stability by allowing neural networks to approximate an $O(1)$ term instead of one that grows explosively. Notably, this technique is equivalent to the $\epsilon$-parameterization introduced in [33, 47]. We denote this method with the suffix +res. After training, new samples are generated via two methods: Reverse SDE and Annealing SDE on manifolds. Although we use "vanilla" to refer to the Iso and Iso+res methods, these two methods also involve manifold information as we utilize the projection operator during the generation process.
To evaluate the quality of the generated samples, we use several metrics to compare the generated distribution with the target distribution. The choice of metrics balances the ability to capture on-manifold distribution differences with reasonable computational cost. We conduct our experiments using the same hyperparameters, except for those specific to each algorithm. Overall, we recommend the Niso+res training algorithm combined with the Reverse SDE generation method, as it performs best in most cases. In experiments on mesh data and the special orthogonal group, we also conduct Riemannian sliced score matching (RSSM, for simplicity) in [8, 17] as a baseline. Further experimental details are provided in Appendix D. We also conduct ablation studies in Appendix E to analyze the bias-variance tradeoff associated with the choice of hyperparameters.
6.1 The hyperplane in 3D space
We first consider a toy model: the hyperplane embedded in a three-dimensional Euclidean space, specifically the set $\mathcal{M}=\{(x,y,z)\in\mathbb{R}^{3}|z=0\}$. The target distribution is a mixture of Gaussian distributions with nine modes located on the plane. In this example, the score function has a closed-form solution that can be explicitly derived.
In Figure 1, we present the error of the tangential component of the learned score function. The results demonstrate that our proposed Niso-DM and Tango-DM significantly reduce the error of the tangential component, as $t$ approaches 0. Table 5 in the Appendix reports the maximum mean discrepancy (MMD), where both Niso-DM and Tango-DM show significant performance improvements.
6.2 Mesh data
We consider the Stanford Bunny [40] and Spot the Cow [7], which contain meshes derived from 3D scanning ceramic figurines of a rabbit and a cow, respectively. To create the target densities, we follow the approach in [32, 19, 6], which utilizes the eigenpairs of the Laplace-Beltrami operator on meshes. The target distribution is chosen as an equal-proportion mixture of density functions corresponding to the 0-th, 500-th, and 1000-th eigenpairs. The manifold $\mathcal{M}$ is a polyhedron composed of triangular faces, differing from the definition provided in Section 2.2. Although the manifold is not smooth along the shared edges between adjacent faces, which form a zero-measure set, numerical performance ensures that this is an acceptable setup.
Table 1 reports the Jensen–Shannon (JS) divergence of the histograms on mesh faces between the generated samples and the original datasets. With the Rescaling technique, both Niso-DM and Tango-DM demonstrate significant improvements; without it, Tango-DM still achieves improvements under the Annealing algorithm. Furthermore, we suspect that the suboptimal performance of RSSM arises from the challenges in achieving a uniform distribution on manifolds with complex geometric structures, such as the narrow neck in "Spot the Cow."
Table 1: Results for the mesh data: JS divergence of the samples generated by the Reverse SDE and Annealing SDE algorithms under different training methods. The symbol "-" indicates that the Reverse SDE sampling method is not applicable to Tango-DM. The results are reported as the mean and standard deviation over five independent runs. The suffix +res indicates the rescaling technique.
Stanford Bunny Spot the Cow
Reversal Annealing Reversal Annealing
[…]
Tango+res - 3.84e-$\pm1$.57e-3 - 3.39e-$\pm1$.74e-3
[…]
6.3 High-dimensional special orthogonal group
We evaluate the performance of our method in a high-dimensional setting on the special orthogonal group $\mathrm{SO}(10)$, a 45-dimensional submanifold embedded in $\mathbb{R}^{100}$. We construct a synthetic dataset drawn from a multimodal distribution on $\mathrm{SO}(10)$, consisting of 5 modes.
Our results demonstrate the effectiveness of the proposed methods. In Table 3, we compare the sliced 1-Wasserstein [2] distances between the generated and the target distributions. These results highlight the accuracy and reliability of our approach in high dimensional manifold settings. Moreover, our approach outperforms RSSM, while the vanilla algorithm does not.
6.4 Alanine dipeptide
We apply our method to alanine dipeptide, a model system frequently examined in computational chemistry. The system’s configuration is characterized by two dihedral angles, $\phi$ and $\psi$ (refer to Figure 4). The manifold consists of the configurations of the system’s 10 non-hydrogen atoms (in $\mathbb{R}^{30}$) with the angle $\phi$ fixed at $70^{\circ}$, which is a level set of the dihedral angle $\phi$.
Table 3 exhibits the results of the 2-Wasserstein distance between the test datasets and the generated samples, showing that our methods outperform the vanilla models under both the Reverse SDE and the Annealing SDE sampling algorithms. However, in this case, training the normalized score functions (+res) does not lead to better results. One possible explanation is that it simultaneously rescales both the tangential and normal components, resulting in the loss of critical tangential information.
Table 2: Results for $\mathrm{SO}(10)$: Sliced 1-Wasserstein distance under different training methods. For more detailed explanations, refer to the caption of Table 1.
Reversal Annealing
[…]
Tango - 1.89e-$\pm2$.05e-3
[…]
Table 3: Results for dipeptide: 2-Wasserstein distance under different training methods. For more detailed explanations, refer to the caption of Table 1.
Reversal Annealing
[…]
7 Related work
Research on manifold-related diffusion models can be broadly categorized into two directions: (1) diffusion models for distributions under the manifold hypothesis, where the underlying manifold is unknown, and (2) diffusion models for distributions on a known manifold, which is the focus of this work. For the former, we discuss the singularity and the scale discrepancy of score functions mentioned in the related work; for the latter, we explore several studies based on known manifolds.
Singularity of score functions.
Several studies on diffusion models under the manifold hypothesis [29, 27, 3] have highlighted the divergence of the score function as $t\to 0$. Beyond empirical observations, recent studies have provided mathematical analysis of diffusion models based on the VPSDE. For instance, Chen et al. [4] present a mathematical analysis under the assumption that each data point lies on a hyperplane. In [29], it is shown that the norm of the score function satisfies $\mathbb{E}\|\nabla_{x}\log p_{t}(x))\|\gtrsim 1/\sqrt{t}$. Furthermore, Lu et al. [27] rigorously prove that this singularity follows a $1/t$ scaling in a strong pointwise sense under the VPSDE, which is similar to our results in (10). Further discussion can be found in [25].
To mitigate this singularity issue, a number of works (e.g. [38]) assume that the initial time is at $t=\epsilon>0$, instead of $t=0$. The study [9] introduces a diffusion model on the product space of position and velocity, employing hybrid score matching to circumvent the singularity. Furthermore, in [37], the score function is parameterized by $s_{\theta}(x,t)=\hat{s}_{\theta}(x)/\sigma_{t}$, where $\hat{s}_{\theta}(x)$ is a neural network.
Scale discrepancy of the score function.
The work [39] uses the singular value decomposition of the score matrix to identify the intrinsic dimension of the manifold, which essentially utilizes the discrepancy of the score function. References [42, 20, 44, 46] investigate the fundamental principles of diffusion models under the manifold hypothesis, by studying the eigen-decomposition of the Jacobian of the score functions. This analysis implicitly relates to the score discrepancy discussed in our work. In particular, the decomposition of the denoiser mapping (Eq. (12)) in [20], which involves a geometry-adaptive harmonic basis, also implies this discrepancy, aligning with (10) in our paper. Papers [1] and [10] study the Memorization and Generalization in Generative Diffusion under the manifold hypothesis through Hidden Manifold Models and Generalized Linear Models. Besides, [16] suggests that a conservative diffusion is guaranteed to yield the correct conclusions when analyzing local features of the data manifold.
While previous works have implicitly touched upon the scale differences in manifold settings, our work explicitly formalizes this phenomenon with rigorous mathematical formulations and introduces novel solutions to address it. Furthermore, we believe our methods can enhance the accuracy of diffusion models during the manifold consolidation phase (as described in [42]) or the collapse transition times (as mentioned in [1, 10]), by refining the computation of the tangential component of the score function.
Additional noise along the normal direction.
Paper [15] proposes adding noise along the normal direction to "inflate" manifolds under the setting of normalizing flows, which is similar to our Niso method. Theorem 4 and Proposition 7 in [15] analyze the relationship between the perturbed distribution and the original distribution on the manifold, which aligns with the conclusions in our equations (12) and (18).
On the other hand, our method differs from [15] in several key aspects. In [15], the added noise has a constant magnitude, requiring additional assumptions (e.g., Q-normally reachable) to prevent interference between noise at different points. In contrast, our analysis (Theorem 3.1) permits the noise magnitude to diminish toward zero, relying on weaker assumptions. Additionally, [15] applies noise inflation prior to training the normalizing flow, whereas our method integrates noise addition directly into the diffusion process, accompanied by the noise dynamics of diffusion models. Finally, while the analysis in [15] focuses on the perturbed density function, our work studies the asymptotic behavior of the score function.
Diffusion models on manifolds
Riemannian Score-based Generative Models [8] extend SDE-based diffusion models to manifold settings by estimating the transition kernels using either the Riemannian logarithmic map or the eigenpairs of the Laplacian–Beltrami operator on manifolds [26]. Riemannian Diffusion Models [17] employ a variational diffusion framework for Riemannian manifolds and, similar to our approach, consider submanifolds embedded in Euclidean space. Additionally, Trivialized Diffusion Models [49] adapt diffusion models from Euclidean spaces to Lie groups.
8 Limitations and future work
Compared to approaches under the manifold hypothesis, our method still requires knowledge of the manifold’s definition. In fields like image and language processing, the manifold structure is often assumed but not explicitly characterized. Future work will focus on extending our approach to scenarios where the manifold is undefined or implicitly represented. Furthermore, Tango-DM is not suitable for the Reverse SDE Sampling algorithm and tends to be slower due to the annealing sampler. Designing Reverse SDE-style algorithms for Tango-DM remains an open direction for future research.
9 Conclusion
We address the multiscale singularity of the score function for manifold-structured data, which limits the accuracy of Euclidean diffusion models. By decomposing the score function into tangential and normal components, we identify the source of scale discrepancies and propose two methods: Niso-DM, which reduces discrepancies with non-isotropic noise, and Tango-DM, which trains only the tangential component. Both methods achieve superior performance on complex manifolds, improving the accuracy and robustness of diffusion-based generative models.
Acknowledgments and Disclosure of Funding
Zichen Liu acknowledges support from the China Scholarship Council (Grant No. 202306010047). Wei Zhang acknowledges support from DFG’s Eigene Stelle (Project No. 524086759). Tiejun Li acknowledges support from the National Key R&D Program of China (Grant No. 2021YFA1003301) and the National Science Foundation of China (Grant No. 12288101).
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: The main claims in the abstract and introduction accurately reflect the paper’s contributions and scope.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: The limitations of the work are discussed in the Appendix.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory assumptions and proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [Yes]
  14. Justification: We provide the assumptions in the statement of the theorem and the proof of the theorem in the Appendix.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental result reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: The code and the datasets are provided.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Justification: We provide open access to the data and the code.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental setting/details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: All the training and test details are provided in the Appendix.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment statistical significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [Yes]
  34. Justification: Error bars are provided in the table.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments compute resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: The information on the computer resources is provided in the Appendix.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code of ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: The research conforms with the NeurIPS Code of Ethics.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [Yes]
  49. Justification: The potential positive societal impacts can be found in the Appendix and there is no negative societal impact of the work.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [N/A]
  54. Justification: This paper poses no such risks.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: The creators or original owners of all assets used in the paper are properly credited.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [N/A]
  64. Justification: The paper does not release new assets.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and research with human subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification: The paper does not involve crowdsourcing nor research with human subjects.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional review board (IRB) approvals or equivalent for research with human subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: The paper does not involve crowdsourcing nor research with human subjects.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
  76. 16.
    Declaration of LLM usage
  77. Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research? Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
  78. Answer: [N/A]
  79. Justification: The LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research.
  80. Guidelines:
    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.
Appendix A Proof of Theorem 3.1
Preparation.
For simplicity of notation, we use $|\cdot|$ to denote the Euclidean norm in the following two sections. We assume that $\tilde{x}\in\mathbb{R}^{n}$ is fixed. Let us recall the expression
$p_{\sigma}(\tilde{x})=\int_{\mathcal{M}}p_{0}(x)p_{\sigma}(\tilde{x}|x)\mathrm{d}\sigma_{\mathcal{M}}(x),\quad\sigma>0\,.$ (23)
Let $x^{*}\in\mathcal{M}$ be the unique minimizer of the minimization problem
$x~|.\min_{x\in\mathcal{M}}$|x-$\tilde${x}|. (24)
In the following, we study the limit of the integral in (23) over $\mathcal{M}$ as $\sigma\rightarrow 0$ by a change of variables using the exponential map on $\mathcal{M}$ at $x^{*}$. To this end, let us first derive some asymptotic expressions related to the exponential map.
Denote by $T_{x^{*}}\mathcal{M}=\{v|\nabla\xi(x^{*})^{T}v=0,v\in\mathbb{R}^{n}\}$ the tangent space of $\mathcal{M}$ at $x^{*}$. Given $v\in T_{x^{*}}\mathcal{M}$, let $\gamma_{v}(t)\in\mathbb{R}^{n}$ be the geodesic curve on $\mathcal{M}$ starting from $\gamma_{v}(0)=x^{*}$ at $t=0$ such that $\dot{\gamma}_{v}(0)=v$. We have Taylor’s expansion
$\gamma_{v}(t)=x^{*}+vt+\frac{1}{2}\ddot{\gamma}_{v}(0)t^{2}+O(t^{3})\,,\quad\mbox{as}~t\rightarrow 0\,.$ (25)
Since $x^{*}=\gamma_{v}(0)$ is the minimizer of (24), we know that $|\gamma_{v}(t)-\tilde{x}|^{2}$ (as a function of $t$) attains its minimum at $t=0$. Therefore, taking the derivative and setting $t=0$, we obtain
$v\cdot(x^{*}-\tilde{x})=0,\quad\forall v\in T_{x^{*}}\mathcal{M}\,.$ (26)
An expression for the second-order derivative $\ddot{\gamma}_{v}(0)$ in (25) is not available in general. However, by differentiating the identity $\big(I-P(\gamma_{v}(t))\big)\dot{\gamma}_{v}(t)=0$ with respect to $t$, and setting $t=0$, we obtain the equation (i.e. the expression for the orthogonal component of $\ddot{\gamma}_{v}(0)$)
$\big(I-P(x^{*})\big)\ddot{\gamma}_{v}(0)=\sum_{j,j^{\prime}}\frac{\partial P_{\cdot j}}{\partial x_{j^{\prime}}}(x^{*})v_{j}v_{j^{\prime}}\,$, (27)
where $\frac{\partial P_{\cdot j}}{\partial x_{j^{\prime}}}(x^{*})$ denotes the vector in $\mathbb{R}^{n}$ whose $i$-th component is $\frac{\partial P_{ij}}{\partial x_{j^{\prime}}}(x^{*})$, for $1\leq i\leq n$.
The exponential map $\exp_{x^{*}}:T_{x^{*}}\mathcal{M}\rightarrow\mathcal{M}$ is well-defined in the neighbourhood $\mathcal{O}$ of $v=0\in T_{x*}\mathcal{M}$, and it is related to the geodesic curves by $\exp_{x^{*}}(v)=\gamma_{\hat{v}}(|v|)$, where $\hat{v}=v/|v|$. Therefore, from (25) we have
$\exp_{x^{*}}(v)=x^{*}+v+\frac{|v|^{2}}{2}\ddot{\gamma}_{\hat{v}}(0)+O(|v|^{3})\,,\quad\mbox{as}~|v|\rightarrow 0\,.$ (28)
Next, let us derive an expression for $|\exp_{x^{*}}(v)-\tilde{x}|^{2}$ when $|v|$ is small. Using (26) and (28), we can obtain
$|\exp_{x^{*}}(v)-\tilde{x}|^{2}=|x^{*}-\tilde{x}|^{2}+|v|^{2}\Big(1+\ddot{\gamma}_{\hat{v}}(0)\cdot(x^{*}-\tilde{x})\Big)+O(|v|^{3})\,,\quad\mbox{as}~|v|\rightarrow 0\,.$ (29)
Noticing that $x^{*}-\tilde{x}$ is orthogonal to the tangent space $T_{x^{*}}\mathcal{M}$ (see (26)), using (27) we can compute the second term on the right hand side of (29)
$|v|^{2}\ddot{\gamma}_{\hat{v}}(0)\cdot(x^{*}-\tilde{x})=|v|^{2}\Big((I-P(x^{*}))\ddot{\gamma}_{\hat{v}}(0)\Big)\cdot(x^{*}-\tilde{x})=\sum_{i,j,j^{\prime}}\frac{\partial P_{ij}}{\partial x_{j^{\prime}}}(x^{*})v_{j}v_{j^{\prime}}(x^{*}_{i}-\tilde{x}_{i})\,.$ (30)
Substituting the above expression into (29) we get
$|\exp_{x^{*}}(v)-\tilde{x}|^{2}=|x^{*}-\tilde{x}|^{2}+v^{T}Sv+O(|v|^{3})\,,\quad\mbox{as}~|v|\rightarrow 0\,$, (31)
where $S\in\mathbb{R}^{n\times n}$ is a matrix whose entries are
$S_{jj^{\prime}}=\delta_{jj^{\prime}}+\frac{1}{2}\sum_{i}\Big(\frac{\partial P_{ij^{\prime}}}{\partial x_{j}}+\frac{\partial P_{ij}}{\partial x_{j^{\prime}}}\Big)(x^{*})(x^{*}_{i}-\tilde{x}_{i})\,,\quad\mbox{for}~1\leq j,j^{\prime}\leq n\,.$ (32)
From the Gershgorin circle theorem [41], all eigenvalues of $S$ are greater than $1-\frac{1}{2}\max_{j}\sum_{i,j^{\prime}}|(\frac{\partial P_{ij^{\prime}}}{\partial x_{j}}+\frac{\partial P_{ij}}{\partial x_{j^{\prime}}})(x^{*})(x^{*}_{i}-\tilde{x}_{i})|$. Furthermore, this lower bound is positive, as $\|x^{*}-\tilde{x}\|_{\infty}<\frac{1}{n^{2}M_{1}}$. Therefore, the matrix $S$ is positive definite. With the above expansions, we prove the two claims.
Proof of the first claim.
We consider the general case where $\tilde{x}\in\mathbb{R}^{n}$ may not belong to $\mathcal{M}$. Since $p_{\sigma}(\tilde{x}|x)$ is the probability density of Gaussian distribution $\mathcal{N}(x,\sigma^{2}I)$, we can derive
$\begin{split}\nabla_{\tilde{x}}\log p_{\sigma}(\tilde{x})=&\frac{1}{p_{\sigma}(\tilde{x})}\int_{\mathcal{M}}p_{0}(x)\frac{x-\tilde{x}}{\sigma^{2}}p_{\sigma}(\tilde{x}|x)\mathrm{d}\sigma_{\mathcal{M}}(x)\\ =&\frac{x^{*}-\tilde{x}}{\sigma^{2}}+\frac{1}{p_{\sigma}(\tilde{x})}\int_{\mathcal{M}}p_{0}(x)\frac{x-x^{*}}{\sigma^{2}}p_{\sigma}(\tilde{x}|x)\mathrm{d}\sigma_{\mathcal{M}}(x)\,.\end{split}$ (33)
On the one hand, since $x^{*}$ is the unique minimizer of (24), there exists $\delta>0$, such that $|x-\tilde{x}|^{2}\geq|x^{*}-\tilde{x}|^{2}+\delta$ for all $x\in\mathcal{M}\setminus\mathcal{O}^{\prime}$, where $\mathcal{O}^{\prime}:=\exp_{x^{*}}(\mathcal{O})$. Therefore, we have
$\Big|\int_{\mathcal{M}\setminus\mathcal{O}^{\prime}}p_{0}(x)\frac{x-x^{*}}{\sigma^{2}}p_{\sigma}(\tilde{x}|x)\mathrm{d}\sigma_{\mathcal{M}}(x)\Big|$ (34)
$\leq$ $(2\pi\sigma^{2})^{-\frac{n}{2}}\bigg[\int_{\mathcal{M}\setminus\mathcal{O}^{\prime}}p_{0}(x)|x-x^{*}|\mathrm{d}\sigma_{\mathcal{M}}(x)\bigg](\sigma^{-2}e^{-\frac{\delta}{2\sigma^{2}}})e^{-\frac{|x^{*}-\tilde{x}|^{2}}{2\sigma^{2}}}$
$=$ $o(e^{-\frac{\delta}{4\sigma^{2}}})e^{-\frac{|x^{*}-\tilde{x}|^{2}}{2\sigma^{2}}}\,.$
On the other hand, using the fact that $\exp_{x^{*}}$ is a diffeomorphism on $\mathcal{O}$ with $|\det D\exp_{x^{*}}(v)|\equiv 1$, applying the change of variables $x=\exp_{x^{*}}(v)$, we get
$\int_{\mathcal{O}^{\prime}}p_{0}(x)\frac{x-x^{*}}{\sigma^{2}}p_{\sigma}(\tilde{x}|x)\mathrm{d}\sigma_{\mathcal{M}}(x)$ (35)
$=$ $(2\pi\sigma^{2})^{-\frac{n}{2}}\int_{\mathcal{O}}p_{0}(\exp_{x^{*}}(v))\frac{\exp_{x^{*}}(v)-x^{*}}{\sigma^{2}}e^{-\frac{|\exp_{x^{*}}(v)-\tilde{x}|^{2}}{2\sigma^{2}}}\mathrm{d}v\,.$
Using the expansion (28), we can write
$p_{0}(\exp_{x^{*}}(v))(\exp_{x^{*}}(v)-x^{*})=p_{0}(x^{*})v+\frac{1}{2}|v|^{2}p_{0}(x^{*})\ddot{\gamma}_{\hat{v}}(0)+(\nabla^{\mathcal{M}}p_{0}(x^{*})\cdot v)v+O(|v|^{3})\,.$ (36)
Hence, using the expansions (31) and (36), we can derive the integral in (35) as
$\int_{\mathcal{O}^{\prime}}p_{0}(x)\frac{x-x^{*}}{\sigma^{2}}p_{\sigma}(\tilde{x}|x)\mathrm{d}\sigma_{\mathcal{M}}(x)$ (37)
$=$ $\int_{\mathcal{O}}\frac{1}{\sigma^{2}}\Big(p_{0}(x^{*})v+(\nabla^{\mathcal{M}}p_{0}(x^{*})\cdot v)v+\frac{1}{2}|v|^{2}p_{0}(x^{*})\ddot{\gamma}_{\hat{v}}(0)+O(|v|^{3})\Big)e^{-\frac{v^{T}Sv+O(|v|^{3})}{2\sigma^{2}}}\mathrm{d}v$
$\quad\quad\cdot(2\pi\sigma^{2})^{-\frac{n}{2}}e^{-\frac{|x^{*}-\tilde{x}|^{2}}{2\sigma^{2}}}$
$=$ $\int_{\mathcal{O}_{\sigma}}\Big[p_{0}(x^{*})\frac{v^{\prime}}{\sigma}+(\nabla^{\mathcal{M}}p_{0}(x^{*})\cdot v^{\prime})v^{\prime}+\frac{1}{2}|v^{\prime}|^{2}p_{0}(x^{*})\ddot{\gamma}_{\hat{v}^{\prime}}(0)+O(\sigma)\Big]e^{-\frac{v^{\prime T}Sv^{\prime}+O(\sigma)}{2}}\mathrm{d}v^{\prime}$
$\quad\quad\cdot(2\pi)^{-\frac{n}{2}}\sigma^{d-n}e^{-\frac{|x^{*}-\tilde{x}|^{2}}{2\sigma^{2}}}$
$=$ $O(1)\sigma^{d-n}e^{-\frac{|x^{*}-\tilde{x}|^{2}}{2\sigma^{2}}}\,$,
where the second equality follows from a change of variables by $v=\sigma v^{\prime}$ and $\mathcal{O}_{\sigma}:=\{\sigma^{-1}v|v\in\mathcal{O}\}$, and the last equality follows from the fact that the integral $\int_{\mathcal{O}_{\sigma}}\frac{v^{\prime}}{\sigma}e^{-\frac{v^{\prime T}Sv^{\prime}+O(\sigma)}{2}}\mathrm{d}v^{\prime}$ is $O(1)$. Combining (34) and (37), we arrive at
$\int_{\mathcal{M}}p_{0}(x)\frac{x-x^{*}}{\sigma^{2}}p_{\sigma}(\tilde{x}|x)\mathrm{d}\sigma_{\mathcal{M}}(x)=O(1)\sigma^{d-n}e^{-\frac{|x^{*}-\tilde{x}|^{2}}{2\sigma^{2}}}\,.$ (38)
Using a similar derivation, for $p_{\sigma}$ we can obtain the expansion
$p_{\sigma}(\tilde{x})=\int_{\mathcal{M}}p_{0}(x)p_{\sigma}(\tilde{x}|x)\mathrm{d}\sigma_{\mathcal{M}}(x)=(2\pi\sigma^{2})^{\frac{d-n}{2}}e^{-\frac{|x^{*}-\tilde{x}|^{2}}{2\sigma^{2}}}\left(p_{0}(x^{*})+O(\sigma)\right)\,.$ (39)
The first claim is obtained by combining (33), (38), and (39).
Proof of the second claim.
Now we consider the case where $\tilde{x}\in\mathcal{M}$. First of all, notice that in this case $S=I$ and (39) simplifies to
$p_{\sigma}(\tilde{x})=(2\pi\sigma^{2})^{\frac{d-n}{2}}(p_{0}(\tilde{x})+O(\sigma))\,.$ (40)
We can derive
[…]
$=$ $-\int_{\mathcal{M}}p_{0}(x)P(x)\nabla_{x}p_{\sigma}(\tilde{x}|x)\mathrm{d}\sigma_{\mathcal{M}}(x)-\int_{\mathcal{M}}p_{0}(x)(I-P(x))\nabla_{x}p_{\sigma}(\tilde{x}|x)\mathrm{d}\sigma_{\mathcal{M}}(x)$
$=$ $-\int_{\mathcal{M}}p_{0}(x)\nabla^{\mathcal{M}}_{x}p_{\sigma}(\tilde{x}|x)\mathrm{d}\sigma_{\mathcal{M}}(x)-\int_{\mathcal{M}}p_{0}(x)(I-P(x))\nabla_{x}p_{\sigma}(\tilde{x}|x)\mathrm{d}\sigma_{\mathcal{M}}(x)$
$=$ $\int_{\mathcal{M}}\nabla^{\mathcal{M}}_{x}p_{0}(x)p_{\sigma}(\tilde{x}|x)\mathrm{d}\sigma_{\mathcal{M}}(x)+\int_{\mathcal{M}}p_{0}(x)(I-P(x))\frac{x-\tilde{x}}{\sigma^{2}}p_{\sigma}(\tilde{x}|x)\mathrm{d}\sigma_{\mathcal{M}}(x)\,$,
where $\nabla^{\mathcal{M}}_{x}$ denotes the gradient operator on $\mathcal{M}$ at $x$ and we used the fact that $\nabla^{\mathcal{M}}_{x}=P(x)\nabla_{x}$ to obtain the third equality, and the last equality follows by using the integration by parts formula on $\mathcal{M}$. For the first term in the last line of (41), similar to (39), we can obtain
$\int_{\mathcal{M}}\nabla^{\mathcal{M}}_{x}p_{0}(x)p_{\sigma}(\tilde{x}|x)\mathrm{d}\sigma_{\mathcal{M}}(x)=(2\pi\sigma^{2})^{\frac{d-n}{2}}(\nabla^{\mathcal{M}}_{x}p_{0}(\tilde{x})+O(\sigma))\,.$ (42)
For the second term in the last line of (41), in analogy to (35), we derive
$\int_{\mathcal{O}^{\prime}}p_{0}(x)(I-P(x))\frac{x-\tilde{x}}{\sigma^{2}}p_{\sigma}(\tilde{x}|x)\mathrm{d}\sigma_{\mathcal{M}}(x)$ (43)
$=$ $(2\pi\sigma^{2})^{-\frac{n}{2}}\int_{\mathcal{O}}p_{0}(\exp_{\tilde{x}}(v))\big(I-P(\exp_{\tilde{x}}(v))\big)\frac{\exp_{\tilde{x}}(v)-\tilde{x}}{\sigma^{2}}e^{-\frac{|\exp_{\tilde{x}}(v)-\tilde{x}|^{2}}{\sigma^{2}}}\mathrm{d}v\,.$
In this case, we have $\tilde{x}\in\mathcal{M}$ and $\tilde{x}=x^{*}$. Therefore, the expansion (31) reduces to $|\exp_{\tilde{x}}(v)-\tilde{x}|^{2}=|v|^{2}+O(|v|^{3})$. Applying (28), we then obtain
[…]
$=$ $\frac{1}{2}(I-P(\tilde{x}))\ddot{\gamma}_{\hat{v}}(0)|v|^{2}-\sum_{j,j^{\prime}}\frac{\partial P_{\cdot j}}{\partial x_{j^{\prime}}}(\tilde{x})v_{j}v_{j^{\prime}}+O(|v|^{3})$
$=$ $-\frac{1}{2}\sum_{j,j^{\prime}}\frac{\partial P_{\cdot j}}{\partial x_{j^{\prime}}}(\tilde{x})v_{j}v_{j^{\prime}}+O(|v|^{3})\,$,
where we have used (26) and (27) to derive the first and the second equality, respectively. We denote $U\in\mathbb{R}^{n\times d}$ the matrix whose columns form an orthonormal basis of $T_{\tilde{x}}\mathcal{M}$. It is straightforward to verify that $U^{T}U=I\in\mathbb{R}^{d\times d}$ and $P(\tilde{x})=UU^{T}$. We can compute (43) as
$\int_{\mathcal{O}^{\prime}}p_{0}(x)(I-P(x))\frac{x-\tilde{x}}{\sigma^{2}}p_{\sigma}(\tilde{x}|x)\mathrm{d}\sigma_{\mathcal{M}}(x)$ (45)
$=$ $(2\pi\sigma^{2})^{-\frac{n}{2}}\int_{\mathcal{O}}(p_{0}(\tilde{x})+O(|v|))\frac{1}{\sigma^{2}}\Big(-\frac{1}{2}\sum_{j,j^{\prime}}\frac{\partial P_{\cdot j}}{\partial x_{j^{\prime}}}(\tilde{x})v_{j}v_{j^{\prime}}+O(|v|^{3})\Big)e^{-\frac{|v|^{2}+O(|v|^{3})}{2\sigma^{2}}}\mathrm{d}v$
$=$ $(2\pi)^{-\frac{n}{2}}\sigma^{d-n}\int_{\mathcal{O}_{\sigma}}(p_{0}(\tilde{x})+O(\sigma))\Big[-\frac{1}{2}\sum_{j,j^{\prime}}\frac{\partial P_{\cdot j}}{\partial x_{j^{\prime}}}(\tilde{x})v^{\prime}_{j}v^{\prime}_{j^{\prime}}+O(\sigma)\Big]e^{-\frac{|v^{\prime}|^{2}+O(\sigma)}{2}}\mathrm{d}v^{\prime}$
$=$ $(2\pi)^{-\frac{n}{2}}\sigma^{d-n}\int_{\mathcal{W}_{\sigma}}\Big[-\frac{1}{2}p_{0}(\tilde{x})\sum_{j,j^{\prime}}\frac{\partial P_{\cdot j}}{\partial x_{j^{\prime}}}(\tilde{x})\Big(\sum_{l=1}^{d}U_{jl}w_{l}\Big)\Big(\sum_{l^{\prime}=1}^{d}U_{j^{\prime}l^{\prime}}w_{l^{\prime}}\Big)$
$\hskip 256.0748pt+O(\sigma)\Big]e^{-\frac{|Uw|^{2}+O(\sigma)}{2}}\mathrm{d}w$
$=$ $(2\pi)^{-\frac{n}{2}}\sigma^{d-n}\int_{\mathcal{W}_{\sigma}}\Big[-\frac{1}{2}p_{0}(\tilde{x})\sum_{j,j^{\prime}}\frac{\partial P_{\cdot j}}{\partial x_{j^{\prime}}}(\tilde{x})(UU^{T})_{jj^{\prime}}+O(\sigma)\Big]e^{-\frac{|w|^{2}+O(\sigma)}{2}}\mathrm{d}w$
$=$ $(2\pi)^{-\frac{n}{2}}\sigma^{d-n}\Big[-\frac{1}{2}p_{0}(\tilde{x})\sum_{j,j^{\prime}}\frac{\partial P_{\cdot j}}{\partial x_{j^{\prime}}}(\tilde{x})P_{jj^{\prime}}(\tilde{x})+O(\sigma)\Big]\,$,
where the third equality follows from a further change of variables with $v^{\prime}=Uw$ and $\mathcal{W}_{\sigma}=\{\sigma^{-1}U^{T}v|v\in\mathcal{O}\}$, the fourth equality follows from $|Uw|^{2}=|w|^{2}$ and the fact that the integral converges to an integral with respect to a standard Gaussian density. Combining (41), (42), and (45), and using the fact that the corresponding integral on $\mathcal{M}\setminus\mathcal{O}^{\prime}$ is $o(e^{-\frac{\delta}{4\sigma^{2}}})$, we derive (11) from
$\nabla_{\tilde{x}}\log p_{\sigma}(\tilde{x})=\frac{\nabla_{\tilde{x}}p_{\sigma}(\tilde{x})}{p_{\sigma}(\tilde{x})}=\nabla^{\mathcal{M}}_{\tilde{x}}\log p_{0}(\tilde{x})-\frac{1}{2}\sum_{j,j^{\prime}}\frac{\partial P_{\cdot j}}{\partial x_{j^{\prime}}}(\tilde{x})P_{jj^{\prime}}(\tilde{x})+O(\sigma).$ (46)
[…]
This implies that $P(\tilde{x})T(\tilde{x})=0$. Therefore, multiplying P(x~)P($\tilde${x}) on both sides of (46) confirms (12). The second claim is obtained.
Appendix B Proof of Theorem 4.1
Preparation.
We first study the properties of $\Sigma_{\sigma}(x)$. Recall that for $x\in\mathcal{M}$,
$\Sigma_{\sigma}(x)=\sigma^{2}I+\sigma^{2\alpha}\nabla\xi(x)\left(\nabla\xi(x)^{T}\nabla\xi(x)\right)^{-1}\nabla\xi(x)^{T}=\sigma^{2}I+\sigma^{2\alpha}N(x)N(x)^{T}.$ (48)
By the Sherman–Morrison–Woodbury formula [31],
$\Sigma_{\sigma}(x)^{-1}=$ $\frac{1}{\sigma^{2}}\left(I-\frac{1}{1+\sigma^{2-2\alpha}}N(x)N(x)^{T}\right)$ (49)
$=$ $\frac{1}{\sigma^{2}}P(x)+\frac{1}{\sigma^{2\alpha}}\frac{1}{1+\sigma^{2-2\alpha}}(I-P(x)).$
Besides,
$\det\Sigma_{\sigma}(x)=\sigma^{2d}(\sigma^{2\alpha})^{n-d}(1+\sigma^{2-2\alpha})^{n-d}.$ (50)
Proof of the first claim.
We first estimate the order of the following quadratic form:
[…]
We apply the same change of variables $x=\exp_{x^{*}}(v)$ as in the proof in Section A and $x=x^{*}+v+O(|v|^{2})$. We obtain
[…]
$=$ $\Big(x^{*}-\tilde{x}+v+O(|v|^{2})\Big)^{T}\left(I-P(x^{*})-\sum_{k}\frac{\partial P(x^{*})}{\partial x_{k}}v_{k}+O(|v|^{2})\right)\Big(x^{*}-\tilde{x}+v+O(|v|^{2})\Big)$
$=$ $|x^{*}-\tilde{x}|^{2}+O(|v|^{2})\,$,
and
($x~)\displaystyle(x-\tilde{x})^{T}P(x$)(x-$\tilde${x}) (53)
$=$ $\Big(x^{*}-\tilde{x}+v+O(|v|^{2})\Big)^{T}\left(P(x^{*})+\sum_{k}\frac{\partial P(x^{*})}{\partial x_{k}}v_{k}+\frac{1}{2}\sum_{k,l}\frac{\partial^{2}P(x^{*})}{\partial x_{k}\partial x_{l}}v_{k}v_{l}+O(|v|^{3})\right)$
$\quad\quad\cdot\Big(x^{*}-\tilde{x}+v+O(|v|^{2})\Big)$
$=$ $|v|^{2}+2\sum_{k,j,j^{\prime}}\frac{\partial P_{jj^{\prime}}}{\partial x_{k}}(x^{*})v_{k}v_{j^{\prime}}(x^{*}_{j}-\tilde{x}_{j})+\frac{1}{2}\sum_{k,l,j,j^{\prime}}\frac{\partial^{2}P(x^{*})}{\partial x_{k}\partial x_{l}}v_{k}v_{l}(x^{*}_{j}-\tilde{x}_{j})(x^{*}_{j^{\prime}}-\tilde{x}_{j^{\prime}})+O(|v|^{3})$
= $v^{T}\tilde{S}v+O(|v|^{3})\,.$
The disappearance of the first-order terms in (52) and (53) is attributed to
($x~)\displaystyle(x^{*}-\tilde{x})^{T}\left(\sum_{k}\frac{\partial P(x^{*})}{\partial x_{k}}v_{k}\right$)($^{*}-\tilde{x}$) (54)
$=$ ($x~)\displaystyle(x^{*}-\tilde{x})^{T}\left(\sum_{k}\frac{\partial(P)^{2}(x^{*})}{\partial x_{k}}v_{k}\right$)($^{*}-\tilde{x}$)
$=$ ($x~)\displaystyle(x^{*}-\tilde{x})^{T}P(x^{*})^{T}\left(\sum_{k}\frac{\partial P(x^{*})}{\partial x_{k}}v_{k}\right)(x^{*}-\tilde{x})+(x^{*}-\tilde{x})^{T}\left(\sum_{k}\frac{\partial P(x^{*})}{\partial x_{k}}v_{k}\right)P(x^{*}$)($^{*}-\tilde{x}$)
$=$ $0\,$,
and $\tilde{S}\in\mathbb{R}^{n\times n}$ is a matrix whose entries are defined by
$\tilde{S}_{jj^{\prime}}=\delta_{jj^{\prime}}+\sum_{i}\left(\frac{\partial P_{ij^{\prime}}}{\partial x_{j}}+\frac{\partial P_{ij}}{\partial x_{j^{\prime}}}\right)(x^{*})(x^{*}_{i}-\tilde{x}_{i})+\frac{1}{2}\sum_{k,l}\frac{\partial^{2}P_{kl}(x^{*})}{\partial x_{j}\partial x_{j^{\prime}}}(x^{*}_{k}-\tilde{x}_{l})(x^{*}_{k}-\tilde{x}_{l}),$ (55)
for $1\leq j,j^{\prime}\leq n$. By the Gershgorin circle theorem [41] and the condition $\|x^{*}-\tilde{x}\|_{\infty}<\min\{1,\frac{2}{n^{2}(4M_{1}+M_{2})}\}$, the matrix $\tilde{S}$ is also positive definite. From (52), (53), and (55), $p_{\sigma}(\tilde{x}|x)$ can be written as
$p_{\sigma}(\tilde{x}|x)=(2\pi)^{-\frac{n}{2}}\sigma^{-d}(\sigma^{-\alpha})^{n-d}(1+o(1))e^{-\frac{1}{2\sigma^{2}}v^{T}\tilde{S}v-\frac{1}{\sigma^{2\alpha}}|x^{*}-\tilde{x}|^{2}+\frac{1}{\sigma^{2}}O(|v|^{3})+\frac{1}{\sigma^{2\alpha}}O(|v|^{2})}.$ (56)
Noting that $(x-\tilde{x})^{T}\Sigma_{\sigma}(x)^{-1}(x-\tilde{x})\geq\frac{1}{1+\sigma^{2-2\alpha}}\frac{1}{\sigma^{2\alpha}}|x-\tilde{x}|^{2}$, there exists $\delta>0$ such that $(x-\tilde{x})^{T}\Sigma_{\sigma}(x)^{-1}(x-\tilde{x})-\frac{1}{\sigma^{2\alpha}}|x^{*}-\tilde{x}|^{2}\geq\frac{\delta}{\sigma^{2\alpha}}$ for all $x\in\mathcal{M}\setminus\mathcal{O}^{\prime}$, where $\mathcal{O}^{\prime}:=\exp_{x^{*}}(\mathcal{O})$. For any $m\in\mathbb{R}$, we have
$\sigma^{m}\int_{\mathcal{M}}|x-x^{*}|p_{0}(x)p_{\sigma}(\tilde{x}|x)\mathrm{d}\sigma_{\mathcal{M}}(x)$ (57)
$\leq$ $\sigma^{m}e^{-\frac{1}{\sigma^{2\alpha}}|x^{*}-\tilde{x}|^{2}-\frac{\delta}{\sigma^{2\alpha}}}\int_{\mathcal{M}}(2\pi)^{-\frac{n}{2}}|\det\Sigma_{\sigma}(x)|^{-\frac{1}{2}}|x-x^{*}|p_{0}(x)\mathrm{d}\sigma_{\mathcal{M}}(x)$
$\leq$ $e^{-\frac{1}{\sigma^{2\alpha}}|x^{*}-\tilde{x}|^{2}}o(e^{-\frac{\delta}{4\sigma^{\alpha}}}).$
Similarly, the order of $\Sigma_{\sigma}(x)^{-1}(\tilde{x}-x)$ can be estimated as
$-\Sigma_{\sigma}(x)^{-1}(\tilde{x}-x)$ (58)
[…]
$=$ $\frac{1}{\sigma^{2}}\left(P(x^{*})+\sum_{k}\frac{\partial P}{\partial x_{k}}(x^{*})v_{k}+O(|v|^{2})\right)\big(x^{*}-\tilde{x}+v+O(|v|^{2})\big)$
$+\frac{1}{\sigma^{2\alpha}}\frac{1}{1+\sigma^{2-2\alpha}}\left(I-P(x^{*})-\sum_{k}\frac{\partial P}{\partial x_{k}}(x^{*})v_{k}+O(|v|^{2})\right)(x^{*}-\tilde{x}+v+O(|v|^{2}))$
$=$ $\frac{1}{\sigma^{2}}(v+Hv+O(|v|^{2}))+\frac{1}{\sigma^{2\alpha}}\frac{1}{1+\sigma^{2-2\alpha}}(x^{*}-\tilde{x}-Hv)+\frac{1}{\sigma^{2\alpha}}O(|v|^{2}),$
where we denote $H\in\mathbb{R}^{n\times n}$ with $H_{ij}=\sum_{l}\frac{\partial P_{il}}{\partial x_{j}}(\tilde{x})(x^{*}_{l}-x_{l})$. Therefore,
[…]
$=$ $\int_{\mathcal{M}}p_{0}(x)\left(-\Sigma_{\sigma}(x)^{-1}(\tilde{x}-x)-\frac{x^{*}-\tilde{x}}{\sigma^{2\alpha}}\frac{1}{1+\sigma^{2-2\alpha}}\right)p_{\sigma}(\tilde{x}|x)\mathrm{d}\sigma_{\mathcal{M}}(x)$
$=$ $\int_{\mathcal{O}}p_{0}(x)\left(\frac{1}{\sigma^{2}}(v+Hv+O(|v|^{2}))-\frac{1}{\sigma^{2\alpha}}\frac{1}{1+\sigma^{2-2\alpha}}Hv+\frac{1}{\sigma^{2\alpha}}O(|v|^{2})\right)p_{\sigma}(\tilde{x}|x)\mathrm{d}\sigma_{\mathcal{M}}(x)$
$+e^{-\frac{1}{\sigma^{2\alpha}}|x^{*}-\tilde{x}|^{2}}o(e^{-\frac{\delta}{4\sigma^{\alpha}}})$
$=$ $\int_{\mathcal{O}}\sigma^{-d}\Big(p_{0}(x^{*})+O(|v|)\Big)\Big(\frac{1}{\sigma^{2}}(v+Hv+O(|v|^{2}))-\frac{1}{\sigma^{2\alpha}}\frac{1}{1+\sigma^{2-2\alpha}}Hv+\frac{1}{\sigma^{2\alpha}}O(|v|^{2})\Big)$
$\quad\quad\cdot e^{-\frac{1}{2\sigma^{2}}v^{T}\tilde{S}v+\frac{1}{\sigma^{2}}O(|v|^{3})+\frac{1}{\sigma^{2\alpha}}O(|v|^{2})}\mathrm{d}v\cdot(2\pi)^{-\frac{n}{2}}(\sigma^{-\alpha})^{n-d}e^{-\frac{1}{\sigma^{2\alpha}}|x^{*}-\tilde{x}|^{2}}(1+o(1))$
$+e^{-\frac{1}{\sigma^{2\alpha}}|x^{*}-\tilde{x}|^{2}}o(e^{-\frac{\delta}{4\sigma^{\alpha}}})$
$=$ $\int_{\mathcal{O}_{\sigma}}\Big(p_{0}(x^{*})+O(\sigma)\Big)\Big(\frac{1}{\sigma}(I+H)v^{\prime}-\frac{\sigma^{1-2\alpha}}{1+\sigma^{2-2\alpha}}Hv^{\prime}+O(1)\Big)e^{-v^{\prime T}S_{1}v^{\prime}+O(\sigma)+O(\sigma^{2-2\alpha})}\mathrm{d}v^{\prime}$
$\quad\quad\cdot(2\pi)^{-\frac{n}{2}}(\sigma^{-\alpha})^{n-d}e^{-\frac{1}{\sigma^{2\alpha}}|x^{*}-\tilde{x}|^{2}}+e^{-\frac{1}{\sigma^{2\alpha}}|x^{*}-\tilde{x}|^{2}}o(e^{-\frac{\delta}{4\sigma^{\alpha}}})$
$=$ $\sigma^{(d-n)\alpha}e^{-\frac{1}{\sigma^{2\alpha}}|x^{*}-\tilde{x}|^{2}}O(\sigma^{(1-2\alpha)\wedge 0}),$
where the third equality follows from a change of variables by $v=\sigma v^{\prime}$ and $\mathcal{O}_{\sigma}:=\{\sigma^{-1}v|v\in\mathcal{O}\}$, and $O(\sigma)+O(\sigma^{2-2\alpha})=O(\sigma^{(2-2\alpha)\wedge 1})$. Besides, using a similar derivation, we can obtain the expansion for $p_{\sigma}$,
$p_{\sigma}(\tilde{x})=\sigma^{(d-n)\alpha}e^{-\frac{1}{\sigma^{2\alpha}}|x^{*}-\tilde{x}|^{2}}\left(p_{0}(x^{*})+O(\sigma)\right).$ (60)
The first claim is obtained by combining (59) and (60).
Proof of the second claim.
Now we consider the case where $\tilde{x}\in\mathcal{M}$. In this case, $\tilde{x}=x^{*}$. By the change of variables $x=\exp_{x^{*}}(v)$, the quadratic form becomes $(x-\tilde{x})^{T}\Sigma_{\sigma}(x)^{-1}(x-\tilde{x})=\frac{1}{\sigma^{2}}(|v|^{2}+O(|v|^{3}))+\frac{1}{\sigma^{2\alpha}}O(|v|^{4})$. Similar to (41) and (45), we aim to show that
$\nabla_{\tilde{x}}p_{\sigma}(\tilde{x})-\int_{\mathcal{M}}\nabla^{\mathcal{M}}_{x}p_{0}(x)p_{\sigma}(\tilde{x}|x)\mathrm{d}\sigma_{\mathcal{M}}(x)$ (61)
$=$ $(2\pi)^{-\frac{n}{2}}\sigma^{(d-n)\alpha}p_{0}(\tilde{x})\Big[-\frac{1}{2}\sum_{j,j^{\prime}=1}^{n}\nabla_{\tilde{x}}P_{jj^{\prime}}(\tilde{x})P_{jj^{\prime}}(\tilde{x})+O(\sigma^{1\wedge(2-2\alpha)})\Big].$
First, by integration by parts,
$\nabla_{\tilde{x}}p_{\sigma}(\tilde{x})-\int_{\mathcal{M}}\nabla^{\mathcal{M}}_{x}p_{0}(x)p_{\sigma}(\tilde{x}|x)\mathrm{d}\sigma_{\mathcal{M}}(x)$ (62)
$=$ $\int_{\mathcal{M}}p_{0}(x)\left(\nabla_{\tilde{x}}p_{\sigma}(\tilde{x}|x)+\nabla^{\mathcal{M}}_{x}p_{\sigma}(\tilde{x}|x)\right)\mathrm{d}\sigma_{\mathcal{M}}(x)$
$=$ $\int_{\mathcal{M}}p_{0}(x)\left[(I-P(x))\Sigma_{\sigma}(x)^{-1}(x-\tilde{x})-\frac{1}{2}\sum_{i,j,k}(x_{i}-\tilde{x}_{i})P_{\cdot k}\frac{\partial(\Sigma_{\sigma}^{ij})^{-1}}{\partial x_{k}}(x)(x_{j}-\tilde{x}_{j})\right]$
$\quad\quad\cdot p_{\sigma}(\tilde{x}|x)\mathrm{d}\sigma_{\mathcal{M}}(x),$
where $\Sigma_{\sigma}^{ij}$ refers to the $(i,j)$ entry of the matrix $\Sigma_{\sigma}$. From (49), we have
$(I-P(x))\Sigma_{\sigma}(x)^{-1}(x-\tilde{x})=\frac{1}{\sigma^{2\alpha}}\frac{1}{1+\sigma^{2-2\alpha}}(I-P(x))(x-\tilde{x})=\frac{1}{\sigma^{2\alpha}}O(|v|^{2})$ (63)
and
$\frac{1}{2}\sum_{i,j,k}(x_{i}-\tilde{x}_{i})P_{\cdot k}\frac{\partial}{\partial x_{k}}(\Sigma_{\sigma}^{ij})^{-1}(x)(x_{j}-\tilde{x}_{j})$ (64)
$=$ $\frac{1}{2\sigma^{2}}\sum_{i,j,k}(x_{i}-\tilde{x}_{i})P_{\cdot k}\frac{\partial}{\partial x_{k}}P_{ij}(x)(x_{j}-\tilde{x}_{j})+\frac{1}{\sigma^{2\alpha}}O(|v|^{2})$
$=$ $\frac{1}{2\sigma^{2}}\sum_{i,j,k}P_{\cdot k}\frac{\partial P_{ij}}{\partial x_{k}}(\tilde{x})v_{i}v_{j}+\frac{1}{2\sigma^{2}}O(|v|^{3})+\frac{1}{\sigma^{2\alpha}}O(|v|^{2}).$
Next, we continue to calculate (62) using (63) and (64). There exists $\delta>0$ and $\mathcal{O}^{\prime}$ such that
$(x-\tilde{x})^{T}\Sigma_{\sigma}(x)^{-1}(x-\tilde{x})\geq\frac{1}{2\sigma^{2\alpha}}\frac{|x-\tilde{x}|^{2}}{1+\sigma^{2-2\alpha}}>\frac{\delta}{\sigma^{2\alpha}},~\forall\sigma>0,~x\in\mathcal{M}\setminus\mathcal{O}^{\prime}.$ (65)
Therefore, the value of the integral (62) in $x\in\mathcal{M}\setminus\mathcal{O}^{\prime}$ is $o(e^{-\frac{\delta}{4\sigma^{\alpha}}})$. Using the same derivation as in (45), we have
$\nabla_{\tilde{x}}p_{\sigma}(\tilde{x})-\int_{\mathcal{M}}\nabla^{\mathcal{M}}_{x}p_{0}(x)p_{\sigma}(\tilde{x}|x)\mathrm{d}\sigma_{\mathcal{M}}(x)$ (66)
$=$ $\int_{\mathcal{O}^{\prime}}p_{0}(\tilde{x}+O(|v|))\left(-\frac{1}{2\sigma^{2}}\sum_{i,j,k}P_{\cdot k}\frac{\partial P_{ij}}{\partial x_{k}}(\tilde{x})v_{i}v_{j}+\frac{1}{2\sigma^{2}}O(|v|^{3})+\frac{1}{\sigma^{2\alpha}}O(|v|^{2})\right)$
$\quad\quad\cdot p_{\sigma}(\tilde{x}|\exp_{x^{*}}(v))\mathrm{d}v+o(e^{-\frac{\delta}{4\sigma^{\alpha}}})$
$=$ $(2\pi)^{-\frac{n}{2}}\sigma^{(d-n)\alpha}p_{0}(\tilde{x})\Big[-\frac{1}{2}\sum_{k,j,j^{\prime}}P_{\cdot k}\frac{\partial P_{jj^{\prime}}}{\partial x_{k}}(\tilde{x})P_{jj^{\prime}}(\tilde{x})+O(\sigma^{1\wedge(2-2\alpha)})\Big]+o(e^{-\frac{\delta}{4\sigma^{\alpha}}}).$
Therefore, (61) holds. Besides,
$\sum_{j,j^{\prime}}\nabla P_{jj^{\prime}}P_{jj^{\prime}}=$ $\sum_{k,j,j^{\prime}}\nabla(P_{jk}P_{kj^{\prime}})P_{jj^{\prime}}$ (67)
$=$ $\sum_{k,j,j^{\prime}}\nabla P_{kj^{\prime}}P_{jk}P_{jj^{\prime}}+\sum_{k,j,j^{\prime}}\nabla P_{jk}P_{kj^{\prime}}P_{jj^{\prime}}$
$=$ $2\sum_{j,j^{\prime}}\nabla P_{jj^{\prime}}P_{jj^{\prime}}$
implies $\sum_{j,j^{\prime}}\nabla P_{jj^{\prime}}(\tilde{x})P_{jj^{\prime}}(\tilde{x})=0$. Therefore, (61) becomes
$\nabla_{\tilde{x}}p_{\sigma}(\tilde{x})-\int_{\mathcal{M}}\nabla^{\mathcal{M}}_{x}p_{0}(x)p_{\sigma}(\tilde{x}|x)\mathrm{d}\sigma_{\mathcal{M}}(x)=(2\pi)^{-\frac{n}{2}}\sigma^{(d-n)\alpha}p_{0}(\tilde{x})O(\sigma^{1\wedge(2-2\alpha)})\,.$ (68)
Similarly, we have
$\int_{\mathcal{M}}\nabla^{\mathcal{M}}_{x}p_{0}(x)p_{\sigma}(\tilde{x}|x)\mathrm{d}\sigma_{\mathcal{M}}(x)=(2\pi)^{-\frac{n}{2}}\sigma^{(d-n)\alpha}\left(\nabla_{\tilde{x}}^{\mathcal{M}}p_{0}(\tilde{x})+O(\sigma^{1\wedge(2-2\alpha)})\right),$ (69)
$p_{\sigma}(\tilde{x})=\int_{\mathcal{M}}p_{0}(x)p_{\sigma}(\tilde{x}|x)\mathrm{d}\sigma_{\mathcal{M}}(x)=(2\pi)^{-\frac{n}{2}}\sigma^{(d-n)\alpha}\left(p_{0}(\tilde{x})+O(\sigma^{1\wedge(2-2\alpha)})\right).$ (70)
Using (68), (69), and (70), the second claim is obtained via
$\nabla_{\tilde{x}}\log p_{\sigma}(\tilde{x})=\frac{\nabla_{\tilde{x}}p_{\sigma}(\tilde{x})}{p_{\sigma}(\tilde{x})}=\nabla_{\tilde{x}}^{\mathcal{M}}\log p_{0}(\tilde{x})+O(\sigma^{1\wedge(2-2\alpha)})\,.$ (71)
Appendix C Algorithm details
Figure 2 illustrates the schematic diagrams of the vanilla method, along with our proposed Niso-DM and tango-DM.
Figure 2: Schematic diagrams of the vanilla method (left), our proposed Niso-DM (middle), and tango-DM (right).
C.1 Details of the Tango loss
We first prove the validity of the loss (20). Note that
$\mathbb{E}_{x_{t}}\|s^{\parallel}_{\theta}(x_{t},t)-P(x_{t})\nabla_{x_{t}}\log p_{\sigma_{t}}(x_{t})\|^{2}$ (72)
$=$ $\mathbb{E}_{x_{t}}\|P(x_{t})s_{\theta}(x_{t},t)\|^{2}-2\mathbb{E}_{x_{t}}\langle P(x_{t})s_{\theta}(x_{t},t),P(x_{t})\nabla_{x_{t}}\log p_{\sigma_{t}}(x_{t})\rangle+C$
$=$ $\mathbb{E}_{x_{t}}\|P(x_{t})s_{\theta}(x_{t},t)\|^{2}-2\int_{\mathbb{R}^{n}}\langle P(x_{t})s_{\theta}(x_{t},t),\nabla_{x_{t}}p_{\sigma_{t}}(x_{t})\rangle\mathrm{d}x_{t}+C$
$=$ $\mathbb{E}_{x_{t}}\|P(x_{t})s_{\theta}(x_{t},t)\|^{2}-2\int_{\mathbb{R}^{n}}\left\langle P(x_{t})s_{\theta}(x_{t},t),\nabla_{x_{t}}\int_{\mathcal{M}}p_{\sigma_{t}}(x_{t}|x)p_{0}(x)\mathrm{d}x\right\rangle\mathrm{d}x_{t}+C$
$=$ $\mathbb{E}_{x_{t}}\|P(x_{t})s_{\theta}(x_{t},t)\|^{2}-2\int_{\mathbb{R}^{n}}\int_{\mathcal{M}}\left\langle P(x_{t})s_{\theta}(x_{t},t),\nabla_{x_{t}}\log p_{\sigma_{t}}(x_{t}|x)\right\rangle p_{\sigma_{t}}(x_{t}|x)p_{0}(x)\mathrm{d}x\mathrm{d}x_{t}+C$
$=$ $\mathbb{E}_{x_{t}}\|P(x_{t})s_{\theta}(x_{t},t)\|^{2}-2\mathbb{E}_{x,x_{t}}\left\langle P(x_{t})s_{\theta}(x_{t},t),\nabla_{x_{t}}\log p_{\sigma_{t}}(x_{t}|x)\right\rangle+C$
$=$ $\mathbb{E}_{x,x_{t}}\|s^{\parallel}_{\theta}(x_{t},t)-P(x_{t})\nabla_{x_{t}}\log p_{\sigma_{t}}(x_{t}|x)\|^{2}+C_{1}\,$,
where $C$ and $C_{1}$ are constants independent of $\theta$. Therefore, the minimizer of the loss (20) satisfies that $s_{\theta}^{\parallel}(x,t)=P(x)\nabla_{x}\log p_{\sigma_{t}}(x)$. The overall loss calculation is summarized in Algorithm 1.
Algorithm 1 The overall loss calculation with the Tango loss
0: neural network $s_{\theta}$, threshold $c_{\mathrm{tango}}$
1: Sample $t\sim\mathcal{U}(0,1)$, $x\sim p_{0}$
2:if $\sigma_{t}\geq c_{\mathrm{tango}}$ then
3:  Calculate loss: $\lambda_{t}\mathbb{E}_{x,x_{t}}\|s_{\theta}(x_{t},t)-\nabla_{x_{t}}\log p_{\sigma_{t}}(x_{t}|x)\|^{2}$
4:else
5:  Calculate loss: $\lambda_{t}\mathbb{E}_{x,x_{t}}\|P(x_{t})s_{\theta}(x_{t},t)-P(x_{t})\nabla_{x_{t}}\log p_{\sigma_{t}}(x_{t}|x)\|^{2}$
6:end if
C.2 Details of the sampling algorithm
Algorithms 2 and 3 provide detailed descriptions of the Reverse SDE and the Annealing SDE on manifolds, respectively. The threshold $\tilde{\sigma}$ in Algorithm 3 is set to $c_{\mathrm{niso}}$ and $c_{\mathrm{tango}}$ for Niso-DM and Tango-DM, respectively.
Algorithm 2 Reverse SDE Solver
0: trained neural network $s_{\theta}$, total number of discrete SDE steps NN, projection operator $\pi$
[…]
3:while $t>0$ do
4:  $z\sim\mathcal{N}(0,I)$
5:  $x_{t}\leftarrow x_{t}+g(t)^{2}s_{\theta}(x_{t},t)\Delta t+g(t)\sqrt{\Delta t}z$
6:  $t\leftarrow t-\Delta t$
7:end while
8: %Projection onto the manifold
[…]
10:return $x_{t}$
Algorithm 3 Annealing SDE on manifolds
0: trained neural network $s_{\theta}$, total number of discrete SDE steps NN, threshold $\tilde{\sigma}$, number of steps $n_{0}$ and step size $\alpha_{\mathrm{ld}}$ for the Langevin dynamics on manifolds, projection operator $\pi$
[…]
2: %Stage 1: Reverse SDE
3:while $\sigma_{t}\geq\tilde{\sigma}$ do
4:  $z\sim\mathcal{N}(0,I)$
5:  $x_{t}\leftarrow x_{t}+g(t)^{2}s_{\theta}(x_{t},t)\Delta t+g(t)\sqrt{\Delta t}z$
6:  $t\leftarrow t-\Delta t$
7:end while
8: %Stage 2: Annealing SDE on manifolds
[…]
10:while $t>0$ do
11:  for $i=0$ to $n_{0}$ do
12:   $z\sim\mathcal{N}(0,I)$, $z_{1}\leftarrow P(x)z$
13:   $x_{t}^{\prime}\leftarrow x_{t}+\alpha_{\mathrm{ld}}P(x_{t})s_{\theta}(x_{t},t)+\sqrt{2\alpha_{\mathrm{ld}}}z_{1}$
14:   $x_{t}\leftarrow\pi(x_{t}^{\prime})$
15:  end for
16:  $t\leftarrow t-\Delta t$
17:end while
18:return $x_{t}$
Appendix D Experimental details
The details of each experiment are provided in the subsections below. The neural networks used are multilayer perceptrons (MLPs) with SiLU activations. Models are trained using PyTorch, utilizing the Adam optimizer with a fixed learning rate, and gradients are clipped when their $2$-norm exceeds a predefined threshold. An exponential moving average of the model weights [30] is applied with a decay rate of $0.999$. In each run, the dataset is divided into training and test sets with a ratio $8:2$. All experiments are conducted on either a single Tesla V100-PCIE-32GB GPU or an NVIDIA A40 GPU with 48 GB of memory. The values of all parameters used in the experiments are listed in Table 4.
When the rescaling technique is not applied, the time reweighting coefficient $\lambda_{t}$ is defined as $\sigma_{t}^{2}$. In contrast, when the rescaling technique is applied, $\lambda_{t}$ is defined as $\sigma_{t}w_{t}$, where $w_{t}$ represents the scaling factor of the optimal score function. Specifically, $w_{t}$ corresponds to $\sigma_{t}$ in Iso-DM, $\sqrt{\sigma_{t}^{2}+c_{\mathrm{niso}}^{2}}$ in Niso-DM, and $\max\{\sigma_{t},c_{\mathrm{tango}}\}$ in Tango-DM.
Table 4: Parameters in our experiments. $\sigma_{\min}$, $\sigma_{\max}$, and $T$ are the parameters of the SDE. $c_{\mathrm{niso}}$ and $c_{\mathrm{tango}}$ are the parameters for Niso-DM and Tango-DM, respectively. NN, $n_{0}$, and $\alpha_{\mathrm{ld}}$ are the hyperparameters of the sampling algorithm. $N_{\mathrm{epoch}}$, $B$, $\mathrm{lr}$, and $\mathrm{clip}$ denote the total epochs, the batch size, the learning rate, and the gradient clipping threshold during training. $N_{\mathrm{node}}$ and $N_{\mathrm{layer}}$ represent the number of nodes per layer and the number of hidden layers in the neural networks, respectively. $\mathcal{D}$ indicates the dataset size.
Parameters Hyperplane Bunny Spot $\mathrm{SO}(10)$ dipeptide
$\sigma_{\min}$ 0.001 0.001 0.001 0.0005 0.0001
$\sigma_{\max}$ 3 3 3 3 5
$T$ 1 1 1 1 1
$c_{\mathrm{niso}}$ 0.2 0.002 0.002 0.01 0.005
$c_{\mathrm{tango}}$ 0.2 0.002 0.002 0.05 0.005
NN 500 200 200 500 500
$n_{0}$ 10 20 10 10 10
$\alpha_{\mathrm{ld}}$ 0.01 0.05 0.05 0.05 0.05
$N_{\mathrm{node}}$ 64 256 256 512 512
$N_{\mathrm{layer}}$ 3 3 3 3 5
$N_{\mathrm{epoch}}$ 200 20000 20000 5000 6000
$B$ 512 4096 4096 512 1024
$\mathrm{lr}$ 0.0005 0.0005 0.0005 0.001 0.0005
$\mathrm{clip}$ 10 10 10 1 10
$\mathcal{D}$ 50000 60000 60000 50000 99999
D.1 Riemannian sliced score matching
Diffusion models constrained to a manifold often rely on complex geometric constructs, such as the heat kernel or logarithmic mapping, to compute transition probabilities. These dependencies limit the efficient computation of the denoising score matching loss. In contrast, Riemannian sliced score matching (RSSM [8, 17]) offers broader applicability to general manifolds. This is achieved by leveraging the implicit score matching loss [8], combined with Hutchinson trace estimation [18]. The corresponding training loss is expressed as:
$\ell_{\text{RSSM}}(t,\theta):=\mathbb{E}_{x_{t}}\|s_{\theta}(x_{t},t)\|^{2}+2\mathbb{E}_{z\sim\mathcal{N}(0,I),x_{t}}\left[z^{T}P(x_{t})^{T}\nabla_{x_{t}}s_{\theta}(x_{t},t)P(x_{t})z\right].$ (73)
D.2 The hyperplane in 3D space
Table 5: Results for the Hyperplane: MMD under different training methods. In this example, the Reverse SDE sampling method is applicable to Tango-DM due to the complete decoupling of tangential and normal components. For more detailed explanations, refer to the caption of Table 1.
Reversal Annealing
[…]
The target distribution is a mixture of Gaussian distributions with nine modes located on the plane. Specifically, the means of the nine modes are (1,1)(-1,-1), (1,0)(-1,0), (1,1)(-1,1), (0,1)(0,-1), $(0,0)$, $(0,1)$, (1,1)(1,-1), $(1,0)$, and $(1,1)$, and each Gaussian has a standard deviation of $0.3$.
D.3 Mesh data
To create the datasets, we first obtain the clamped eigenfunctions of the Laplacian operator on a mesh that has been upsampled threefold for the original mesh. The target distribution is chosen as an equal-proportion mixture of density functions corresponding to the 0-th, 500-th, and 1000-th eigenpairs. For the distribution on the cow mesh, we additionally discarded the points on the horns and tail.
For $x\notin\mathcal{M}$, the closest point $\pi(x)$ on a triangular mesh surface is determined as follows: First, $x$ is projected onto the planes of all triangular faces in the mesh using the face normals and vertex positions. Barycentric coordinates are then computed to check whether the projected points lie inside the triangles. If a point falls outside a triangle, it is further projected onto the nearest edge to ensure it remains on the triangle’s boundary. Subsequently, the Euclidean distance between the original query point $x$ and all projected points is calculated, and the closest point $\pi(x)$ is selected by identifying the minimum distance.
We assign the normal direction $n(\pi(x))$ of $\pi(x)$ as the normal direction for $x$ ($x\notin\mathcal{M}$). When $\pi(x)$ is located inside a triangle of the mesh, the normal of the triangle is chosen as the normal vector at $\pi(x)$ (i.e. $n(x)=n(\pi(x))$). However, when $\pi(x)$ lies on an edge (or vertex) of the mesh, it is simultaneously associated with two (or three) triangles. In this case, we select the normal vector of the triangle with the smallest index as the normal direction for $\pi(x)$. Here, we leverage the property of the torch.argmin function.
D.4 High-dimensional special orthogonal group
The manifold $\mathrm{SO}(10)$ is defined as $\{Q\in\mathbb{R}^{10\times 10}|QQ^{T}=I_{10},\operatorname{det}(Q)=1\}$, which represents (a connected component of) the zero-level set of the map $\xi:\mathbb{R}^{100}\rightarrow\mathbb{R}^{55}$. The components of $\xi$ correspond to the upper triangular portion of the matrix $QQ^{T}-I_{10}$. The dataset is constructed as a mixture of 5 wrapped normal distributions. Each wrapped normal distribution is the image (under the exponential map) of a normal distribution defined in the tangent space at a pre-selected center. For more details, refer to [23]. In addition, Figure 3 shows that compared with Iso-DM, the samples from Niso-DM have a better match with the ground truth.
Figure 3: Results for $\mathrm{SO}(10)$. Empirical densities of the statistics $\operatorname{tr}(S),\operatorname{tr}(S^{2}),\operatorname{tr}(S^{4})$, and $\operatorname{tr}(S^{5})$ for Iso-DM (red), Niso-DM (purple) and the ground truth (green). Samples from Niso-DM (purple) better match the ground truth (green).
D.5 Alanine dipeptide
To generate the dataset, we follow [23] to obtain the target distribution on the manifold $\mathcal{M}=\{x\in\mathbb{R}^{30}|\phi(x)=-70^{\circ}\}$. We also ensure that the generated distribution is $\mathrm{SE}(3)$-invariant (i.e., invariant under rotations and translations) by incorporating alignment before feeding the data into the neural network.
Figure 4: Illustration of the Alanine dipeptide system: The dihedral angles $\phi$ and $\psi$ are defined by atoms whose indices are $5,7,9,15$ and $7,9,15,17$, respectively. This figure is from [23].
Appendix E Ablation Study
Recall that the noise scale $\sigma_{t}$ is chosen as $\sigma_{t}=\sigma_{\min}(\sigma_{\max}/\sigma_{\min})^{t/T}$ and $0<\sigma_{\min}\ll 1$. When $t$ is sufficiently small, $\sigma_{t}$ approaches $\sigma_{\min}$, and the singularity of the score function primarily depends on $\sigma_{\min}$. As $\sigma_{\min}$ decreases, the singularity becomes pronounced.
For Iso-DM, there exists a bias-variance trade-off in the choice of $\sigma_{\min}$. A very small $\sigma_{\min}$ reduces bias but increases variance, resulting in a larger overall error. In this case, $\nabla_{x}\log p_{\sigma_{\min}}(x)$ approximates $\nabla_{x}\log p_{0}(x)$ well; however, severe multiscale discrepancies between its tangential and normal components hinder learning. Conversely, a very large $\sigma_{\min}$ increases bias and reduces variance, yet the overall error still increases. In this scenario, $\nabla_{x}\log p_{\sigma_{\min}}(x)$ fails to serve as a good approximation of $\nabla_{x}\log p_{0}(x)$, although the singularity issue is less pronounced. Figure 5(a) illustrates this phenomenon, where the red line represents the error for Iso-DM.
For Niso-DM, the singularity is governed by the additional noise scale $c_{\mathrm{niso}}$, rather than $\sigma_{\min}$. As a result, our method exhibits greater robustness as $\sigma_{\min}$ decreases. The blue line in Figure 5(a) demonstrates the stability of the error for Niso-DM under varying $\sigma_{\min}$.
Figures 5(b) and 5(c) show the impact of $c_{\mathrm{niso}}$ in Niso-DM and $c_{\mathrm{tango}}$ in Tango-DM on the error. The selection of these hyperparameters also involves a bias-variance trade-off.
(a) Error vs $\sigma_{\min}$
(b) Error vs $c_{\mathrm{niso}}$
(c) Error vs $c_{\mathrm{tango}}$
Figure 5: Ablation Studies for $\mathrm{SO}(10)$: The solid line denotes the mean, while the shaded area indicates the standard deviation. (a) The error of the distribution generated by the Reverse SDE algorithm under the Iso-DM (red) and Niso-DM (blue) methods with varying $\sigma_{\min}$. (b) The impact of $c_{\mathrm{niso}}$ on the error with $\sigma_{\min}=0.001$. (c) The impact of $c_{\mathrm{tango}}$ on the error with $\sigma_{\min}=0.001$.
In Tables 6-9, we present the numerical results of ablation studies on $\sigma_{\min}$, $c_{\mathrm{niso}}$, and $c_{\mathrm{tango}}$, for the $\mathrm{SO}(10)$ experiment. We report computational errors under different parameter settings to assess the sensitivity of our method to hyperparameter choices. The results demonstrate that reductions in hyperparameter values may not substantially affect the performance of our method.
Table 6: The impact of $\sigma_{\min}$ on the error under the Niso algorithm.
$\sigma_{\min}$ 1e-5 1e-4 5e-4 1e-3 1e-2
[…]
Table 7: The impact of $c_{\mathrm{niso}}$ on the error under the Niso algorithm.
$c_{\mathrm{niso}}$ 1e-3 5e-3 1e-2 5e-2
[…]
Table 8: The impact of $\sigma_{\min}$ on the error under the Tango algorithm.
$\sigma_{\min}$ 1e-5 1e-4 5e-4 1e-3 1e-2
[…]
Table 9: The impact of $c_{\mathrm{tango}}$ on the error under the Tango algorithm.
$c_{\mathrm{tango}}$ 5e-3 1e-2 5e-2 1e-1
[…]
Appendix F Discussion
F.1 Computational cost of Tango-DM
While Tango-DM tends to be slower due to the annealing sampler, the runtime is primarily determined by the number of steps of the inner Langevin dynamics. For example, in the toy experiment R2inR3, the sampling time for Reverse SDE is 1.68 seconds, while for Annealing SDE, the sampling time increases to 5.62 seconds with an inner step of 5 and to 9.40 seconds with an inner step of 10. The first phase of Annealing SDE (see Algorithm 3) requires only 0.82 seconds. Overall, the computational cost of Annealing SDE is approximately 3–6 times higher than that of Reverse SDE. These measurements, conducted on a standard laptop, highlight the relative computational expense.
F.2 Learned manifold case
Unlike previous works on distributions on a known manifold, our method avoids relying on geometric information like geodesics or heat kernels, greatly reducing computational complexity. However, compared to approaches under the manifold hypothesis, our method still requires knowledge of the manifold’s definition, including projection operators. In fields like image and language processing, the manifold structure is often assumed but not explicitly known.
Next, we discuss how to extend our method to learned manifolds, using the AutoEncoder as an example. Specifically, we first train an Encoder $\phi_{\theta_{1}}:\mathcal{M}\to\mathbb{R}^{d}$ and a Decoder $\psi_{\theta_{2}}:\mathbb{R}^{d}\to\mathcal{M}$, which provide a parameterized representation of the manifold. The subspace spanned by $\nabla\psi_{\theta_{2}}(\phi_{\theta_{1}}(x))$ corresponds to the tangent space $T_{x}\mathcal{M}$ at point $x$ on the manifold. Once the basis of the tangent space is obtained, the score function can be further decomposed into its tangential and normal components, which enables the implementation of our proposed methods. Intuitively, the success of this approach hinges on the Autoencoder’s ability to accurately capture the underlying manifold structure. In this work, we leave this approach as a direction for future research.
F.3 Assumptions in Theorem 3.1 and Theorem 4.1
The conclusions in Theorem 3.1 and Theorem 4.1 hold pointwise on the manifold, so in the proof, we only require local boundedness. Specifically, the following assumption is sufficient:
  • For any $x\in\mathcal{M}$, there exists $\delta>0$, such that $\sup_{y\in\mathcal{M}_{x}^{\delta}}\max_{1\leq i,j,j^{\prime}\leq n}\left|\frac{\partial P_{ij}}{\partial y_{j^{\prime}}}(y)\right|<+\infty$, where $\mathcal{M}_{x}^{\delta}=\{\arg\min_{x^{*}\in\mathcal{M}}|\tilde{x}-x^{*}|^{2}\mid|\tilde{x}-x|<\delta\}$.
To make the theorem statement more concise, we adopted a stronger assumption:
  • $\sup_{x\in\mathcal{M}}\max_{1\leq i,j,j^{\prime}\leq n}\left|\frac{\partial P_{ij}}{\partial x_{j^{\prime}}}(x)\right|<+\infty$.
For compact manifolds, the uniform boundedness assumptions always hold, which implies that the local boundedness assumptions hold. Similarly, the boundedness assumption in Theorem 4.1 can also be relaxed to local boundedness.
Appendix G Impact statement
This paper contributes to the advancement of generative models for data with manifold structures, providing a deeper understanding of the singularity of the score function. Specifically, we identify the scale discrepancies between the tangential and normal components of the score function, which sheds light on key challenges in modeling data on manifolds. We believe that our work bridges the gap between generative models specifically designed for manifolds and those aimed at handling data under manifold assumptions. While this study may have broader societal implications, none require particular emphasis at this stage.